### v4

# Reputation in the Shadow of Exit

Abstract
I study reputation formation in repeated games where player actions endogenously determine the probability the game permanently ends. Permanent exit can render reputation useless even to a patient long-lived player whose actions are perfectly monitored, in stark contrast to canonical commitment payoff theorems. However, I identify tight conditions for the long-run player to attain their Stackelberg payoff in the unique Markovian equilibria. Along the way, I highlight the role of Markovian strategies in pinning down the value of reputation formation. I apply my results to repeated global games with exit, and obtain new predictions about regime survival and a dynamic foundation for risk dominant selection.
Keywords: Repeated Global Games, Reputation, Regime Change, Endogenous Exit.
JEL Codes: C73, D74, I14.
1 Introduction
Economists have long understood the extent to which reputation effects could refine payoffs in complete information repeated games as players become arbitrarily patient. Famously, the seminal work of Fudenberg and Levine (1989) showed a patient long-run player could attain a payoff arbitrarily close to their Stackelberg payoff in the presence of reputational incentives, in stark contrast to folk theorems which could not give a payoff prediction in the absence of those incentives. Of crucial importance to this result is the arbitrary patience assumption: the long-run player expects with high probability to play arbitrarily many periods of the game. This then allows Fudenberg and Levine (1989) to apply their following observation to prove their Stackelberg payoff bound result:
If the long-run player chooses his Stackelberg strategy in every period, there is an upper bound on the number of times the short-run players can fail to play as Stackelberg followers.
This simple observation has proven quite powerful; for very general games which persist with (exogenous) arbitrarily high probability for many periods, much work has been done generalizing this observation to arbitrary monitoring structures and stage game payoffs (see Fudenberg and Levine (1992), Gossner (2011) and Luo and Wolitzky (2025)).
There are, however, many settings where the probability the game itself persists depends on the actions of short-run players. Consider for example the chain-store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982). A crucial assumption in their setup is that the incumbent is able to repeatedly maintain their position, even if they unsuccessfully deter entry. Yet in many markets, successful entry by a challenger can displace an incumbent, changing the structure of the game in future periods or forcing the incumbent themselves to exit. Similarly, in the product choice game of Mailath and Samuelson (2006) (Chapter 15.1) and Pei (2020), it may be that a long enough string of low quality forces the firm to exit. Finally, in repeated models of regime change (see Huang (2017)), a successful currency attack could end the current regime (or lead to the toppling of the central bank). Players, realizing this, may change their behavior, which could in turn affect reputational dynamics.
In such games, the expected number of periods play persists may be significantly shorter than the “upper bound on the number of times the short-run players can fail to play as Stackelberg followers,” at which point it is not clear whether it is still possible for the long-run player to leverage their reputational incentives to secure a high payoff. Indeed, absent strong assumptions on the exit technology itself, it is impossible for economists to make any predictions in general about payoffs, restoring the indeterminancy of the folk theorem even with reputational incentives (see Proposition 4). Despite this, the literature has until now ignored the effect exit may have on the long-run player’s incentives to secure reputation; it has been universally assumed that the game will be exogenously infinitely repeated, so that the logic of the Fudenberg and Levine (1989) upper bound can be applied.
This paper studies reputation formation where a long-run player (choosing $c$) interacts with many short-run players (choosing $a$) who can be either a rational type or a committed type playing the same action. The key modeling innovation is that there is an exit function $f(a,c)$ in each period that determines whether the game continues into the next stage, or whether the game permanently ends. In several important cases studied by the literature—the chain store game, the product choice (trust) game, regime change games—the specific exit assumptions can distort canonical payoffs predictions in a few ways. First, there is no meaningful theory of normalized discounted theory of payoffs with endogenous exit. Lemma 1 shows that so long as the exit probability is nontrivial (i.e. never $1$), then any normalized discount stream converges to 0 in the patient limit. Not to be deterred, this paper resolves this problem by dealing directly with the unnormalized payoff streams. This lends itself to new challenges as well—in particular, Proposition 1 shows that the optimal set of commitment postures varies with the discount rate (albeit, in a monotone way). This makes it unclear whether there can exist a “single” commitment posture that remains fixed as $\delta\to 1$, as is often done to establish payoff predictions in the reputation literature. Instead, it necessitates restrictions on the sets of feasible commitment postures, which are handled in the setup to the main result. Thus, even without studying the payoff set of the repeated game, exit already requires significant modification to the standard analysis.
The main result of this paper is that, in the payoff setting studied by this paper described above, two additional conditions are jointly necessary and sufficient for reputation effects to select the long-run player’s Stackelberg payoff as the unique equilibrium (Markovian) payoff:
  1. (1)
    An incentive compatibility condition: the effective discount factor must be such that the long-run player must be willing to take the Stackelberg action when they are maximally rewarded for doing so.
  2. (2)
    A credible belief condition: whenever the long-run player has incentives for reputation formation, short-run players must believe that they will pursue them.
The incentive compatibility condition is captured by an inequality that is assumed for the statement of Theorem 1 and ensures the Stackelberg payoff is attainable at some equilibrium in the (complete information) repeated game (see also Proposition 2). The credible belief condition is captured by a Markovian refinement, that short-run player actions are measurable with respect to their belief about the long-run players’ type. This ensures that short-run players act on their beliefs about the long-run player’s rationality, and rules out “revelation equilibria” that have the long-run player reveal rationality today, but which guarantee exit today with high probability. Proposition 4 constructs revelation equilibria of this form which guarantee the long-run player an arbitrarily low payoff absent the Markovian refinement, even when payoffs satisfy my incentive compatibility condition.
Theorem 1 also leads to novel comparisons of Markovian equilibria. In the complete information game, Markovian equilibrium cannot sustain the Stackelberg payoff even when it is an equilibrium of the repeated game in general (Propositions 2 and 3). However, when the regime faces reputational incentives, the unique Markovian equilibrium is the one where the long-run player plays their Stackelberg action at every period and perfectly pools with the committed type [^2] This contrasts with past work, which shows that even under perfect monitoring there can be a continuum of equilibria, see Pei and Li (2021). Thus, the Markovian refinement not only helps select the regime’s best equilibrium payoff, but allows the regime to do strictly outperform, in the presence of reputational incentives, their best Markovian payoff in the complete information game. With perfect monitoring, this phenomena is new, and is a consequence of exit.
Finally, I apply my model to repeated globalized regime change games with exit. A long-run player endogenously determines (at a cost) some coordination friction, after which short-run players see noisy signals about the regime’s strength and decide to attack. The game continues if and only if sufficiently few players attack that the regime is able to overcome the revolution in that period. Such a setup is similar to the game in Angeletos et al. (2007), but with reputational effects and endogenous regime actions. I use Theorem 1 to establish an asymptotic selection result [^3] The asymptotics here require both the idiosyncratic and aggregate noise to vanish, and the regime to become arbitrarily patient in order to satisfy my incentive compatibility condition., Theorem 2: In the unique Markovian equilibrium, the regime attains their Stackelberg payoff if and only if the average strength of the regime is above the endogenously chosen risk-dominant level. This result gives a new dynamic foundation for risk-dominant selection, even when the state is almost known. Conversely, if the average strength is below the risk dominant level, than the regime dies almost surely in the first period and attains an arbitrarily low payoff. Since observed equilibrium play jumps at the risk-dominant threshold, my result sheds light on why seemingly ex-ante similar regimes can differ dramatically ex-post in both (1) observed cruelty and (2) length of survival.
The remainder of this paper is organized as follows. I next review the related literature. Section 2 introduces the model and some basic properties. Section 3 analyzes the Stackelberg action under endogenous exit. Section 4 establishes a folk theorem and a reputation payoff bound in Markov strategies, the main results of the paper. Section 5 applies the framework to study repeated global games of regime change. Section 6 concludes. Omitted proofs are collected in the appendix.
1.1 Related Literature
This paper contributes to the (extensive) literature on when reputational incentives can select payoffs, started by Kreps and Wilson (1982) and Milgrom and Roberts (1982). I adopt the general framework of the seminal work due to Fudenberg and Levine (1989) and Fudenberg and Levine (1992), who establish the generality of reputational effects in a wide variety of infinitely repeated games with long-run and short-run players (see Fudenberg et al. (1990) and Fudenberg and Levine (1994)). There is a large and rich literature making variations to their model; see Part IV of Mailath and Samuelson (2006) or Mertens et al. (2015) for a survey.
This paper is particularly related to Pei (2024) and Pei (2025), who study endogenous records as a key friction to reputation building. Importantly, they show endogenous records implies there are different roles to the exit probability and patience of the players, and that how these two forces interact can lead to different predictions. I undertake a similar decomposition, but endogenize the exit probability. This paper also relates to Pei and Li (2021), who characterize equilibrium play in the Fudenberg and Levine (1989) model without exit; I characterize unique Markovian play with exit.
Relatedly, Abreu and Gul (2000) and the literature on reputational bargaining (see Fanning and Wolitzky (2022) for a review) assume the game ends when one player concedes, a particular form of exit. While the intuition is similar, the models are quite different: for example, bargaining models are zero-sum, assume long-run long-run players, and have very special exit structures. That these models are economically different can be seen in the new technical challenges that arise in my model and the distinctiveness of the qualitative conclusions that attain.
My paper also relates to models of “bad reputation” (Ely and Valimaki (2003) and Ely et al. (2008)) where a shutdown of statistical learning is similar to exit. I contribute to this literature by first directly working with the exit probability as a primitive instead of a derived object due to monitoring, and show that under this interpretation there are cases where reputation can once again be “good” and benefit the long-run player.
Using Markovian refinements to make predictions in reputational repeated games is not new, and I see my refinement as contributing to a long and historical strand of literature that makes this assumption in order to say something meaningful about equilibrium play. For example, Pei (2016) and Lukyanov et al. (2025) use a Markovian refinement to establish “inverse U-shaped” behavior in labor and goods markets, respectively, to reputational incentives. Giannitsarou and Toxvaerd (2003) use Markovian refinements to render the repeated global game (which is a leading application of my framework) tractable and make predictions. Finally, Bhaskar et al. (2013) provide a foundation for Markov perfection in long-run short-run games with bounded memory: they are the only purifiable equilibrium. In games with reputation, Sannikov and Faingold (2011) show that continuous-time games select a unique sequential equilibrium which is Markovian in the belief of the long-run player’s rationality—exactly my criteria. My paper gives an alternative justification for Markovian criteria, as a minimal way of selecting payoffs in the repeated game with endogenous exit.
Finally, the regime change application builds on the rich literature on equilibrium selection in global games, started by Carlsson and van Damme (1993) and Morris and Shin (1998) (see Morris and Shin (2003) for a survey). Our key innovation in this area is to (1) allow the regime to choose, at a cost, the optimal punishment level, and (2) embed the game into a repeated regime change game reminiscent of Angeletos et al. (2007) and Chassang (2010) but with an i.i.d. state. The second change allows me to analyze reputational effects, and unlike Angeletos et al. (2007) and Angeletos and Pavan (2013), establish the existence of a unique equilibrium. Moreover, the dynamic embedding gives a novel dynamic criteria for risk dominant selection, complementary to the speed of learning interpretation in Koh et al. (2025). The first variation relates to a large literature in global games where some “large” play can take an action that modifies the global game and leads to new applied insights about regime change (see Angeletos et al. (2006), Edmond (2013), and Morris and Shadmehr (2024)). Proposition 1 makes an independent contribution in this direction, by establishing a novel comparative static that rationalizes why stronger or more patient regimes are more likely to be cruel, even when doing so harms the regime.
2 Model
2.1 The Stage Game
There are $N+1$ players, $i\in[N]\cup\{L\}$, representing a long-run player $L$ who interacts in each period with $N$ short-run players. Each short-run player $i$ chooses an action from a compact set $a_{i}\in A_{i}$; let $a=\{a_{i}\}\in\prod_{i}A_{i}=A$ denote a profile of short-run player strategies. The long-run player chooses an action $c\in C\subset\mathbb{R}$, where $C$ is a countable (potentially finite) set including 0444In the applications, the set of actions will be a compact interval. Note this is essentially without loss of generality given Assumption 1: If $f:Q\to\mathbb{R}$ is continuous, where $Q\subset X$ is countable and dense, there is a unique continuous function $\bar{f}:X\to\mathbb{R}$ such that $\bar{f}|_{Q}=f$. However, the countability assumption helps significantly when defining equilibrium beliefs off-path, so I maintain it throughout the general analysis.. Short-run player actions generate finite signal $Y$, drawn independently according to a distribution $\rho(\cdot|a)$ [^5] Short-run player monitoring is irrelevant to the analysis and included only for generality.. The long-run player’s action is perfectly monitored.
The action profile $(a,c)$ jointly determines a survival probability, denoted by a function $f:A\times C\to[0,1]$. This can be interpreted as the survival probability of a regime given a short-run action profile $a$ and long-run action $c$.
The long-run player’s expected stage payoff is thus given by $f(a,c)v(a,c)$, where $v(a,c)$ is the payoff the long-run player attains conditional on surviving. Short-run player $i$ has payoffs $u_{i}(a,c)$. I suppose all payoffs are bounded and continuous in all arguments, and moreover normalize long-run payoffs to be nonnegative: $v(a,c)\geq 0$ for all $(a,c)\in A\times C$. One can interpret this normalizing as a requirement that the regime always prefers to live to fight another day: exit is weakly dominated.
For any long-run player action $c$, denote by $\mathcal{N}(c)$ the Nash correspondence at $c$; that is,
$\mathcal{N}(c)=\left\{a:a_{i}\in\operatorname*{\arg\!\max}_{\bar{a}_{i}\in A_{i}}u_{i}(\bar{a}_{i},a_{-i},c)\right\}.$
I make the following assumptions on the survival probability $f$.
Assumption 1.
$f(a,c)$ is continuous with codomain $(0,1)$. Moreover, $f^{*}(c)=\min_{a\in\mathcal{N}(c)}f(a,c)$ is nondecreasing in $c$.
Continuity is a technical assumption that guarantees the existence of a Stackelberg action for each discount rate. That all survival probabilities are interior can be seen as simply stating that there is always some (potentially arbitrarily small) probability that the regime can survive (or die) regardless of what actions are taken [^6] This helps ensure the likelihood ratio is well-defined. Finally, the requirement $f^{*}(c)$ is nondecreasing in $c$ allows one to interpret the long-run player’s choice $c$ as some action which determines a “probability of exit,” ordered so that larger choices of $c$ imply higher probabilities of survival. Throughout, define
$a^{*}(c)=\min_{a\in\mathcal{N}(c)}\left\{f(a,c)v(a,c)\right\}$
as the set of actions that minimize the long-run player’s stage-game payoff, subject to the fact short-run players are best-replying to each other given $c$ is the long-run player’s action. When there is no loss of ambiguity, I also abuse notation and use $a^{*}(c)$ to mean a selection from the correspondence (for example, $f(a^{*}(c),c)$).
Throughout the analysis, it will be useful (expositionally) to impose the assumption a (myopic) long-run player who need not worry about exit has a dominant action, labeled at $c=0$, and moreover that they would prefer to “milk” their reputation at this dominant action whenever possible.
Assumption 2.
For all $a$, $0=\operatorname*{\arg\!\max}_{c\in C}f(a,c)v(a,c)=\operatorname*{\arg\!\max}_{c\in C}v(a,c)$. For any $c>0$, $v(a^{*}(c),0)>v(a^{*}(0),0)$.
The first part of the assumption implies that the set of payoffs attainable in the one-shot repeated game is exactly $\{f(a,0)v(a,0)\}_{a\in\mathcal{N}(0)}$, and simplifies exposition when constructing punishments for deviating from prescribed equilibrium (though, at the cost of substantially more notational overhead, is not necessary). Let $\bar{V}_{0}(\delta)=\max_{a\in\mathcal{N}(0)}\frac{f(a,0)v(a,0)}{1-\delta f(a,0)}$ denote the long-run player’s largest payoff from playing the static optimum repeatedly, and analogously let $\underline{V}_{0}(\delta)$ be her smallest such payoff. The second part of the assumption guarantees the long-run player strictly benefits from the ability to build a reputation when compared to their worst stage-game Nash payoff. $a^{*}(0)$ can thus be interpreted as the “worst-case” action for the principal, i.e. “exit as often as possible.”
2.2 The Repeated Game
The long-run player plays the stage game in each of infinitely many periods ($t=0,1,2,\dots$), discounting the future at some constant rate $\delta\in(0,1)$. In each period, they face a population of short-run players who take myopic best replies. Short-run player actions are monitored according to $y\in\rho(\cdot|a)\in\Delta(Y)$ for a finite set [^7] Short-run monitoring is irrelevant for the analysis and included only for generality., while long-run player actions are perfectly monitored. Time-$t$ histories are elements $h^{t}=(y_{s},c_{s})_{s=0}^{t-1}$ of past short-run player (and long-run player) actions. Let $H^{t}$ denote the set of time-$t$ histories, $H=\bigcup_{t}H_{t}$ the set of all histories, and $H^{\infty}=(Y\times C)^{\infty}$ the set of all infinite histories. Short-run player strategies are functions $\sigma_{i}:H\to\Delta(A_{i})$, while the long-run player’s strategy is a function $\sigma_{L}:H\to\Delta(C)$ [^8] Following Mertens et al. (2015), there is a canonical mapping from strategies $\sigma:H\to\Delta(\mathcal{X})$ and strategies $\sigma:H^{\infty}\to\Delta(\mathcal{X})$ for $\mathcal{X}\in\{C,A\}$. When it is useful (i.e. when topologizing $\sigma$), I will use these two codomains interchangeably..
Given a sequence of realized actions $(a_{t},c_{t})_{t=0}^{\infty}$, the long-run player’s payoff is
$\mathcal{V}(\{a_{t},c_{t}\})=\sum_{t=0}^{\infty}\left(\delta^{t}v(a_{t},c_{t})\prod_{s=0}^{t}f(a_{s},c_{s})\right).$
Define the long-run player’s lower Stackelberg strategy as
$V_{*}(\delta)=\sup_{c\in C}\inf_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}\delta^{t}v(a,c)$
and their upper Stackelberg strategy as
$V^{*}(\delta)=\sup_{c\in C}\max_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}\delta^{t}v(a,c).$
Importantly, we allow the long-run player’s Stackelberg payoffs (and with it, their optimal commitment action) to change with their discount rate. This is because different commitment actions not only affect stage-game payoffs, but also the probability of reaching each terminal node, and hence more patient agents may choose different commitment actions. This feature is special to games with exit and is absent in traditional repeated games models.
Note that the probability of exit in each period is cumulative in the long-run player’s payoff: once they leave the game, their payoff is zero in all future periods regardless of any hypothetical actions that would have been taken. This distinction is a important consequence of exit in my model; in particular, it implies that even as $\delta\to 1$, actions taken in the first few periods still loom large and can have first order impact.
The long-run player can either be a rational type, $\omega_{R}\in\Omega$, who seeks to maximize their discounted payoff, or one of potentially many committed types, $\omega_{c}\in\Omega$, who plays $c$ at every possible history. Short-run players hold a full-support prior belief $\mu_{0}\in\Delta(\Omega)$ over the long-run player’s type.
For any strategy $\sigma$, define an associated belief system $\mu:H\to\Delta(\Omega)$ which outputs some belief about the long-run players’ type for all histories. Together, $(\sigma,\mu)$ is an assessment. Given an assessment, let $\bar{\sigma}_{L}(h^{t})$ denote the unconditional strategy of the long-run player averaging over $\mu(h^{t})$, and $\sigma_{L}(h^{t};\omega)$ denote the strategy of a type $\omega$ long-run player. Analogously, let $\overline{\mathbb{P}}^{\sigma}$ denote the probability measure over histories generated by the unconditional strategy $\sigma$, and $\mathbb{P}^{\sigma}$ the measure generated when $\omega=\omega_{R}$. The main results of this paper study the set of (rational) long-run player payoffs that can be attained in sequential equilibria $(\sigma,\mu)$, defined as follows:
Definition 1.
An assessment $(\sigma,\mu)$ is a sequential equilibrium if:
  1. (1)
    Sequential Rationality. At every history $h^{t}$, the continuation strategy $\sigma_{L}(\cdot|h^{t})$ maximizes the long-run player’s continuation payoff given $\{\sigma_{i}(h^{t})\}$ and $\mu(\cdot|h^{t})$.
  2. (2)
    Short-Run Best Replies. At every history $h^{t}$, $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(\bar{\sigma}_{L}(h^{t}))$.
  3. (3)
    Bayesian Updating. At every history $h^{t}\in\text{supp}(\overline{\mathbb{P}}^{\sigma})$, $\mu(h^{t})=\mu(\cdot|h^{t},\sigma)$, i.e. the conditional belief over $\Omega$ given history $h^{t}$ and strategy $\sigma$.
  4. (4)
    Consistency. There exists a sequence $(\sigma^{n},\mu^{n})$, with $(\sigma^{n},\mu^{n})\to(\sigma,\mu)$ such that
    1. (a)
      $\text{supp}(\overline{\mathbb{P}}^{\sigma^{n}})=H$ for all $n$, and $\mu(\cdot|h^{t},\sigma^{n})=\mu^{n}(h^{t})\in\Delta(\Omega)^{o}$, and
    2. (b)
      $\sigma_{L}^{n}(\cdot,\omega_{c})=c$ at every history $h^{t}$.
An assessment $(\sigma,\mu)$ is a Markov perfect equilibrium if it is a sequential equilibrium and moreover at any histories $h^{t},h^{s}$,
$\mu(h^{t})=\mu(h^{s})\implies\sigma(h^{t})=\sigma(h^{s})$
Because $\sigma:H\to\Delta(C)$ maps an infinite set into another, the notion of convergence is not immediately clear. Say $(\sigma^{n},\mu^{n})$ converges to $(\sigma,\mu)$ if $\sigma^{n}\to\sigma$ converges in the strong topology (see Lemma 3 of Luo and Wolitzky (2025)) and $\mu^{n}\to\mu$ converges weakly. At any history $h^{t}$ where $\mu(h^{t})$ is degenerate, sequential equilibria are equivalent to subgame perfection. Consistency of equilibrium profiles will be useful when evaluating the value of deviations the long-run player may consider which are off the path of play (and not supported by some commitment type); absent some consistency-type condition, the value of such deviations would be hard to define in a principled way. Note the “trembles” defined here for consistency do not change the prior belief over $\Omega$ (which would then change $\mu^{n}$). A standard compactness argument shows this is without loss of generality (I could add the requirement $\mu_{0}^{n}\to\mu_{0}$ as well), but for simplicity I omit this additional notation.
Markov perfection requires the stronger assumption that at any two histories where short-run player posterior beliefs about long-run play are the same, players take the same actions [^9] It may seem unnatural to require the long-run player take the same action when short-run player beliefs are the same. The definition of Markov perfection can be relaxed to require only that short-run players strategies are measurable in $\mu$; however, since there is a single long-run player, their strategy can then be taken to be (without loss of generality) also measurable in $\mu$, less a public randomization device.. While such a refinement may seem a-priori restrictive, I show it arises naturally when studying the effect exit has on reputational incentives; see Theorem 1.
For a fixed discount rate $\delta$, let $\underline{U}(\delta)$ and $\bar{U}(\delta)$ be the long-run player’s lowest and highest payoff from any sequential equilibrium in the game. [^10] There is some potential ambiguity about whether sequential equilibrium need to exist when $C$ is infinite (see Myerson and Reny (2020) for some discussion and general fixes). However, given $\Omega$ is finite and $C$ is countable, the problem is substantially simpler—I show in Proposition 2 and Theorem 1 that sequential equilibria as defined above exist, so in particular $\underline{U},\bar{U}$ are finite. Similarly, let $\underline{U}^{M}(\delta)$ and $\bar{U}^{M}(\delta)$ be the long-run player’s lowest and highest payoff from any MPE in the game.
2.3 Non-normalized payoffs
It will be useful to discuss in somewhat more detail the effect of non-normalized payoff streams on players discounted payoffs, and to highlight why normalization is infeasible as a way to computing average payoffs. Intuitively, without exit, the stream of normalized discounted payoffs converges to the average payoff as $\delta\to 1$. However, when the game has endogenous exit, so long as exit occurs with high enough probability, the average stream of payoffs will be zero for “most” periods (i.e. if the expected exit time is finite), so normalized payoffs will vanish as players get arbitrarily patient. The lemma formalizes this observation.
Lemma 1.
Let $\{f_{t}\}\subset[0,1]$ be a sequence of exit probabilities. If $\liminf_{t\to\infty}f_{t}<1$, then
$\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=0}^{\infty}\delta^{t}v_{t}\prod_{s=0}^{\infty}f_{s}=0.$
for any feasible payoffs $v_{t}$.
The proof can be found in the Proof of Lemma 1 Lemma 1 implies that it is not fruitful to compare discounted average payoffs in the patient limit, as they will often be zero even when payoffs are meaningfully different.
There are two potential ways to ameliorate this problem. The first is to normalize by the effective discount rate, $f(a,c)\delta$, in each period. However, such a normalization is infeasible and difficult to interpret: $f(a,c)$ is an endogenous object, and thus one should naturally expect that it vary the long-run players’ payoffs. Moreover, except in special cases, the stream of exit probabilities $\{f(a_{t},c_{t})\}_{t=0}^{\infty}$ need not be time-invariant, making it hard to understand what the “average effective discount rate” should be. The second way is to simply work directly with the undiscounted payoffs $\mathcal{V}(\{a_{t},c_{t}\})$; I take this approach in this paper.
This choice, however, leads itself to a novel technical challenge: without normalizing payoffs by the discount rate, deviations today may remain first order (in particular, they are not order $(1-\delta)$ compared to an order $\delta$ variation in the continuation payoff, and are instead order $1$ versus order $\delta$ variation in the continuation payoff). This makes it much harder to discipline the long-run player from deviating to play non-Stackelberg actions, since deviations from potential equilibrium profiles will always remain tempting. Theorem 1 gives a simple economic condition under which the long-run player can successfully establish a (Markovian) reputation, despite this hurdle.
3 Commitment Postures
Let’s start with a basic question intrinsic to our model with exit: How does the discount rate affect the long-run player’s Stackelberg payoff? Define
$c^{*}(\delta)=\operatorname*{\arg\!\max}_{c\in C}\min_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}\delta^{t}f(a,c)^{t+1}v(a,c)$
to be the set of (lower) Stackelberg actions for the long-run player when they discount the future at rate $\delta$. Note the long-run player trades off between a higher survival probability at the (potential) cost of a lower stage-game payoff in each period. My first result characterizes exactly how the long-run player resolves this tradeoff as they get more patient: regardless of the speed at which $v(a,c)$ might decline in $c$, they will always favor more aggressive actions that increase their survival probability as they get more patient.
Proposition 1 (Patience Effect).
$c^{*}(\delta)$ is nonempty, upper hemicontinuous, and increasing (in the strong set order) in $\delta$.
Proof.
We first claim $f^{*}(c)$ is continuous, since for any sequence $c_{n}\to c$,
$\lim\limits_{n\to\infty}f^{*}(c_{n})=\lim\limits_{n\to\infty}\frac{v(a^{*}(c_{n}),c_{n})f(a^{*}(c_{n}),c_{n})}{1-\delta f(a^{*}(c_{n}),c_{n})}=\frac{v(a^{*}(c),c)f(a^{*}(c),c)}{1-\delta f(a^{*}(c),c)}.$
The first inequality is by definition, and the second by upper hemi-continuity of the Nash correspondence. Given continuity of $f^{*}(c)$, Berge’s theorem of the maximum applied to the optimizing set $c^{*}(\delta)=\max_{c\in C}f^{*}(c)$ implies $c^{*}(\delta)$ exists and is upper hemi-continuous. Monotonicity now relies on the following lemma, whose proof is deferred to the Proof of Lemma 2  [^11] Such a lemma appears elementary, but to the best of my knowledge there is not a prior reference. Note the lemma establishes a general comparative statics result for discounted streams of payoffs that cross zero once from below.
Lemma 2.
Let $\{b_{t}\}_{t\in\mathbb{N}}\subset\mathbb{R}$ be a bounded sequence of real numbers, not identically zero, such that $b_{t}<0$ if and only if $t<T$ for some $T\in\mathbb{Z}_{+}$. Then for any $\delta_{1}<\delta_{2}$,
$\sum_{t=0}^{\infty}\delta_{1}^{t}b_{t}\geq 0\implies\sum_{t=0}^{\infty}\delta_{2}^{t}b_{t}>0.$
Fix any $\delta$ and $c^{\prime}>c$; if $f^{*}(c)=0$ for all $c\in C$ then the result is trivial; otherwise the solution to $c^{*}(\delta)$ will always lie in $\{c:f^{*}(c)=0\}$. Thus, it is without loss to consider $(c^{\prime},c)$ where $f^{*}(c^{\prime}),f^{*}(c)>0$. Consider the difference
$\sum_{t=0}^{\infty}\delta^{t}[v(a^{*}(c^{\prime}),c^{\prime})f^{*}(c^{\prime})^{t+1}-v(a^{*}(c),c)f^{*}(c)^{t+1}]$
where I abuse notation to and have $a^{*}(\cdot)$ be any selection from the minimizing correspondence. Note that if $b_{t}$ is defined as
$b_{t}=v(a^{*}(c^{\prime}),c^{\prime})f^{*}(c^{\prime})^{t+1}-v(a^{*}(c),c)f^{*}(c)^{t+1}$
then
$b_{t}\geq 0\iff\frac{v(a^{*}(c^{\prime}),c^{\prime})}{v(a^{*}(c),c)}\geq\left(\frac{f^{*}(c)}{f^{*}(c^{\prime})}\right)^{t+1}.$
The left hand side is strictly positive, while the right hand side is positive, in $(0,1)$ (by Assumption 1), and strictly decreasing in $t$. Thus, there exists a unique $T$ (possibly 0) at which the inequality holds for all $t\geq T$. Lemma 2 then implies, for any $c^{\prime}>c$, $\sum_{t=0}^{\infty}\delta^{t}b_{t}$ is increasing in $\delta$, and hence the objective function $\sum_{t=0}^{\infty}\delta^{t}v(a^{*}(c),c)$ is single-crossing in $(\delta,c)$. The result then follows by the Milgrom and Shannon (1994) comparative statics theorem. ∎
Proposition 1 applies regardless of the speed or curvature of $v$ (in particular, it holds even if $v$ declines very fast relative to $c$), though this comes at the cost of a weak comparative static (every selection from $c^{*}(\delta)$ is nondecreasing in $c$). Note Assumption 2 is not used in the proof. Moreover, it has economic implications in the context of the applications—a regime seeking to fend off a coordinated attack will invest more in survival when they are more patient; longer-lived firms care more about ensures entrants are deterred; patient motorcyclists have a greater incentive to take “separating” actions that prove they are not commitment types, in order to deter exit.
That $c^{*}(\delta)$ varies with the discount rate produces a new order of limits challenge in trying to prove some version of a commitment payoff theorem. In particular, since the long-run player’s Stackelberg action depends on the discount rate, it is not clear that there is a finite set $\Omega$ such that $\Omega\cap c^{*}(\delta)$ is nonempty for all $\delta$ close to $1$. Consequently, for any fixed set of commitment types, it may be infeasible to approximate a patient committed sender.
To obviate this problem, I index and simplify the set of commitment postures, $\Omega(\delta)=\{\omega_{R},\omega_{c^{*}(\delta)}\}$ (where I abuse notation and use $c^{*}(\delta)$ to refer to a single selection from the set of commitment types). Moreover, as before, I set $\underline{U}(\delta)$ and $\bar{U}(\delta)$ as the infimum and supremum of the long-run player’s best and worst sequential equilibrium payoff, though now I implicitly fix the set of commitment types $\Omega=\Omega(\delta)$, though I still suppress dependence on the specific sequence of prior beliefs $\mu_{0}(\delta)\in\Delta(\Omega(\delta))$ (which itself my also change in $\delta$). Give this notation, the limit infimum (supremum) $\liminf_{\delta\to 1}\underline{U}(\delta)$ (resp. $\limsup_{\delta\to 1}$) should be interpreted as the limit infimum (supremum) of the sequence of lower equilibrium payoffs in the repeated game with discount rate $\delta$, commitment types $\Omega(\delta)$, and some fixed sequence of prior beliefs $\mu_{0}(\delta)\in\Delta(\Omega(\delta))^{o}$. This interpretation implies that statements of the form $\limsup_{\delta\to 1}|V_{*}(\delta)-\underline{U}(\delta)|=0$ can be read as “The long-run player attains almost their Stackelberg payoff so long as there is positive probability players believe they are committed to their Stackelberg action”.
4 Commitment Payoff Theorems
4.1 Complete Information
Consider first the complete information game where there is no uncertainty about the long-run player’s rationality. Recall that a folk theorem holds in the game if, as $\delta\to 1$ the set of payoffs attainable as part of some equilibrium in the complete-information game with discount rate $\delta$ converges to the set of feasible and individually rational payoffs for the long-run player (here the set $V_{0}(\delta),V^{*}(\delta)]$), supposing short-run players’ myopically best reply. Proposition 2 characterizes exactly when a folk theorem holds in our game. Arguments from Fudenberg and Levine (1994), Fudenberg et al. (1994) (which would immediately imply a folk theorem) do not necessarily apply because exit can “loom large,” and thus lead to periods where the effective discount rate is small even as $\delta\to 1$. Thus, there exist games where not all feasible, individually rational payoffs are attainable. The next result characterizes, as a function of long-run player payoffs, exactly when a folk-like theorem holds as a function of the long-run player’s payoff, and formalizes the sense in which efficient payoffs can be sustained only when the future “looms large” given the exit structure. Throughout, let $L(a,c)=\frac{f(a,c)}{f(a,0)}$ be the likelihood ratio of exit when the long-run player chooses $c$ at some action instead of 0 given short-run player action $a$.
Proposition 2.
Let $\mu_{0}(\omega_{R})=1$ and let $(a^{*},c^{*})$ attain $V^{*}(\delta)$. Then
$\lim\limits_{\delta\to 1}[\underline{U}(\delta),\bar{U}(\delta)]=\lim\limits_{\delta\to 1}[\underline{V}_{0}(\delta),V^{*}(\delta)]$
if and only if there exists $\bar{\delta}<1$ such that for all $\delta>\bar{\delta}$ and any $(a^{*},c^{*})$ attaining $V^{*}(\delta)$,
$\delta[V^{*}(\delta)L(a^{*},c^{*})-\underline{V}_{0}(\delta)]\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*}).$
If either of the two (equivalent) statements above holds, say the game admits a folk theorem. Proposition 2 gives a good baseline for when all feasible payoffs can be obtained in some equilibrium; moreover, the relevant incentive compatibility conditions will give, under some monotonicity and uniqueness conditions, an exact characterization of when the long-run player exactly secures their Stackelberg payoff. Moreover, the converse direction of the result implies the long-run player’s best equilibrium payoff is bounded away from their Stackelberg payoff even as players get arbitrarily patient (i.e. the folk theorem fails) so long as the incentive compatibility condition fails.
The proof is relatively straightforward and can be found in the Proof of Proposition 2 The key step in the proof is to write the incentive constraint for the long-run player, supposing they obtain $V^{*}(\delta)$ if they play $c^{*}$ and $\underline{V}_{0}(\delta)$ otherwise and algebraically rearrange terms to conclude it is equivalent to the condition in the proposition. This then implies playing $c^{*}(\delta)$ is self-enforcing if (and only if) the condition in the result is satisfied.
Importantly, though, the equilibria constructed in the proof of Proposition 2 to sustain a high payoff for the long-run player require that punishments off the equilibrium path are non-Markovian, in the sense that they are not measurable in $\mu_{t}(h^{t})$ (which is constant in the complete information case). Given this observation, it is perhaps unsurprisingly that the Markovian restriction can strongly refine the set of payoffs: in the complete information game, it does so exactly by holding the long-run player down to their exact static payoff.
Proposition 3.
Fix any $\delta$. Then $\underline{U}^{M}(\delta),\bar{U}^{M}(\delta)]=[\underline{V}_{0}(\delta),\bar{V}(\delta)]$.
This property of the complete-information Markov game will be important in disciplining the value of deviations in the game with reputational concerns. The proof follows immediately once it becomes clear that the Markovian refinement requires the continuation strategies to be independent of the history (since $\mu(h^{t})=1$ at all histories in all sequential equilibria). For completeness, I give a formal argument in the Proof of Proposition 3 . Note that Proposition 3 applies to any repeated game, not necessarily just those with exit.
4.2 Guaranteeing Stackelberg Payoffs
What classes of reputation results might one hope for in the game with exit? Recall that classical reputation results hold only after many121212“Many” here depends formally on the speed of learning and the prior belief of rationality; see Gossner (2011) for a more formal discussion. (finite) periods have passed, as the long-run player must teach their opponents that they are playing the commitment action. Yet in the game with exit, they may not have the luxury of time; the game may have ended far before short-run player beliefs about long-run play have concentrated on $c^{*}$, keeping the long-run player from securing a high payoff.
Thus exit perturbs the long-run player’s strategic incentives in two ways: first, by making the future less valuable (and hence the promise of future payoffs less motivating to discipline signaling behavior today), and second, by shortening the length of time they have to convince short-run player beliefs about their (committed) play.
The first problem, as evidenced by Proposition 2, can be resolved by adding an appropriate incentive compatibility constraint in $(\delta,L(\cdot,\cdot))$. The second problem, however, is not so easy to fix: indeed, I will need the full force of the Markovian refinement in order to iron out beliefs which are adversarial to guaranteeing the long-run player their Stackelberg payoff. The technical contribution of Theorem 1 is that these two ingredients exactly characterize when one can obtain a commitment payoff theorem in the game with exit.
Definition 2.
Say the long-run player asymptotically satisfies (weak) $c^{*}$ incentive compatibility if there exists $\bar{\delta}<1$ such that for all $\delta>\bar{\delta}$ and each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$\delta|V_{*}(\delta)L(a,c^{*}(\delta))-\bar{V}_{0}(\delta)|>(\geq)\left|v(a,0)-v(a,c^{*}(\delta))L(a,c^{*}(\delta))\right|$
This incentive compatibility condition is stronger than the one found in the folk theorem, Proposition 2. Intuitively, it requires that for all potential short-run player best replies $a$ that can arise, the future looms sufficiently large that playing $c^{*}$ is enforceable against playing 0 when the long-run player can secure their Stackelberg payoff from $c^{*}$ and static Nash otherwise. In the leading application, where the likelihood ratio and stage-game payoffs depend only on $c$ (i.e. $L(a,c^{*})=L(c^{*})$ and $v(a,c)=v(c)$), this condition boils down exactly to the incentive compatibility condition in Proposition 2 (see Corollary 1). For now, it delivers the following theorem, (modulo one last definition for the converse).
Definition 3.
Primitives $(f,v,u)$ are monotone if $A$ can be ordered such that for each fixed $c$, $f(a,c)$ is strictly decreasing in $A$, $v(a,c)$ is nonincreasing, and $\mathcal{N}(\alpha c)$ is decreasing (in the strong set order) in $\alpha$.
Theorem 1.
Let $\Omega(\delta)=\{0,c^{*}(\delta)\}$. If the long-run player asymptotically satisfies $c^{*}$ incentive compatibility, then for all $\mu_{0}$ with $\mu_{0}(\omega_{R})<1$,
$\limsup_{\delta\to 1}|\underline{U}^{M}(\delta)-V_{*}(\delta)|=0$
for any $\mu_{0}(\omega_{R})<1$. Conversely, if $(f,v,u)$ is monotone and the long-run player does not asymptotically satisfy weak $c^{*}$ incentive compatibility, then for all $\delta<1$,
$\limsup_{(\delta,\mu_{0}(\omega_{R}))\to(1,1)}|\underline{U}^{M}(\delta)-V_{*}(\delta)|>0.$
Proof.
The forward direction. Suppose the long-run player asymptotically satisfies $c^{*}$ incentive compatibility, and fix $\delta>\bar{\delta}$ in the definition. We show in all Markov perfect equilibria the long-run player plays their commitment action $c^{*}\in c^{*}(\delta)\cap\Omega$ and short-run players best respond, securing the long-run player $V_{*}(\delta)$. To do so, we establish a number of lemmas whose (technical) proofs are relegated to the Proof of Lemma 3
First, we must discipline the inferences and payoffs that short-run players have when the long-run player does not imitate the commitment type.
Lemma 3.
Let $(\sigma,\mu)$ be a sequential equilibrium. For all $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$, $\mu(h^{t})(\omega_{R})=1$.
Lemma 3 (along with Proposition 3) guarantees that, off the equilibrium path, the highest payoff the long-run player can obtain is $\bar{V}_{0}(\delta)$, regardless of their discount rate. Our next lemma disciplines the set of actions the long-run player can take in any equilibrium.
Lemma 4.
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $\text{supp}(\sigma_{L}(h^{t}))\in\{0,c^{*}(\delta)\}$.
Lemma 4, along with Lemma 3 are enough to imply (following an argument similar to the first part of the proof of Proposition 2 and taking $\alpha=0$), that always playing $c^{*}$ is a Markov perfect equilibrium when $\delta>\bar{\delta}$.
Our last step is to show that this is the only Markov perfect equilibrium. Clearly, always playing 0 cannot be Markov perfect under assumption. To see why, take $\alpha=0$ and fix any $a\in\mathcal{N}(0)$. Rearranging the incentive compatibility constraint implies
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]<f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
so the long-run player can strictly profitably deviate by always choosing $c^{*}$ in any period they were supposed to deterministically play 0. Finally, we show that the long-run player prefers not to mix in Markovian equilibria.
Lemma 5.
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $|\text{supp}(\sigma_{L}(h^{t}))|=1$.
The intuition is that in order to sustain mixing, the long-run player must be indifferent between their upper Nash and lower Stackelberg payoffs, modulo the cost today. When this difference is large and exit does not shut the game down early (the second condition), then this can never be case, which ensures the long-run player is never indifferent between actions.
The proof of the converse is constructive and can be found in the Proof of Theorem 1 (Converse) The basic idea is that, at any $\alpha$ where the incentive compatibility constraint holds with equality, we find an equilibrium where the long-run player mixes at the time $t=0$ public history. Monotonicity of the game then ensures that at mixing histories, they underperform their commitment payoff. The condition that $\mu_{0}$ is small rules out equilibria where $\mu_{0}(\omega_{c^{*}})\approx 1$ and the long-run player can secure a very high payoff in the first period even if they play 0. ∎
The Markovian refinement here disciplines short-run player learning; long-run players, in a Markovian equilibrium, begin teaching short-run players about play at time $t=0$ (instead of needing to slowly form their reputation). Reputation thus only has a “backwards looking” component: instead of needing to teach short-run players of future play by taking some commitment action today, they need only maintain their current reputation that they will continue to do what they have already done. In reputational games with exit, this backwards-looking force is the economically important one (i.e. the one that is not blocked by the exit structure).
Monotonicity of the game is satisfied in natural applications, including the coordination game later studied in this paper. It is satisfied for example whenever the exit probability increases with “more aggressive’ actions $a\in A$, $v$ depends only on the long-run player’s action, and short-run players play a strictly submodular game with unique best replies (as is the case in the globalized coordination game analyzed later) in $(a,c)$ so that higher levels of long-run player actions lead to less aggressive sets of actions. Monotonicity is sufficient to imply mixed Stackelberg strategies cannot do strictly better than pure ones, which rules out the long-run player from benefiting from the mixed Nash equilibria that can arise when Markovian equilibrium need not be pure (i.e. asymptotic $c^{*}$ incentive compatibility is violated, as per Lemma 5).
In general, the incentive compatibility condition given in the theorem can be somewhat unwieldy to interpret. However, it turns out that there is a special subclass of problems in which the ability to attain the Stackelberg payoff (uniquely) is equivalent to asking if it possible to support that payoff in the complete-information game at all.
Corollary 1.
Suppose $v(a,c)=v(c)$, $L(a,c)$ is constant in $a$, $|\mathcal{N}(0)|=1$, and $V^{*}(\delta)=V_{*}(\delta)$. Then the following are equivalent.
  1. (1)
    There exists $\bar{\mu}_{0}>0$ and $\delta<\bar{\delta}$ such that for all $\mu_{0}<\bar{\mu}_{0}$, $\delta>\bar{\delta}$, $\underline{U}^{M}(\delta)=V^{*}(\delta)$.
  2. (2)
    The game admits a folk theorem.
The proof simply verifies that, under the stated conditions, the incentive compatibility conditions of Proposition 2 and Theorem 1 coincide. It is delegated to the Proof of Corollary 1
Theorem 1 and Corollary 1 are the main theoretical takeaways of this paper. First, they identify a sufficient condition—asymptotic $c^{*}$-incentive compatibility—that negates the effect of exit and ensures the long-run player their Stackelberg payoff in Markovian equilibrium. Moreover, by Proposition 3, the long-run player attains a strictly higher payoff when they have incentives for reputation formation than their best Markovian payoff in the game without reputation. Finally, under a monotonicity condition, these conditions are tight—the long-run player attains their Stackelberg payoff exactly when they satisfy $c^{*}$ incentive compatibility. Corollary 1 then implies that, under some regularity conditions, Markovian equilibria uniquely select the long-run player’s payoff only when such a payoff is feasible, highlighting exit as the key friction hindering reputational incentives in this setting.
4.3 Non-Markovian equilibrium
Theorem 1 focuses on Markovian strategies and leaves open the question of when reputation can help refine payoffs away from the Markovian case. Without putting additional restrictions on the survival probability function $f$, is it possible to say anything about when the long-run player can establish a reputation outside of Markovian strategies? I next show the answer is no—even with asymptotic $c^{*}$ incentive compatibility, one cannot refine the payoff set in general, as the long-run player’s payoff may be arbitrarily low.
Proposition 4.
Fix payoffs $(u,v)$ such that there exists $c$ where $\min_{a\in\mathcal{N}(c)}v(a,c)>0$. Then for any $\varepsilon>0$, there exists survival probability function $f$ such that
  1. (1)
    The folk theorem holds.
  2. (2)
    $\limsup_{\mu_{0}(\omega_{R})\to 1,\delta\to 1}\underline{U}_{0}(\delta)\leq\varepsilon$.
The first assumption is a nontriviality condition guaranteeing the existence of a non-Nash commitment posture, conditional on a high probability of survival. The first limit in the second statement guarantees that the probability of commitment is not “so high” that what the long-run player does will not affect their payoffs (i.e. if $\mu_{0}(\omega_{c^{*}})\approx 1$). The second limit guarantees that the folk theorem holds, and shows that when it does reputation effects cannot, in general, refine the set of payoffs.
Proposition 4 highlights the two important roles of Markovian equilibrium in establishing the positive reputation result of Theorem 1. First, it disciplines the feasible payoff set after rationality is revealed (thus decreasing incentives for “revealing rationality”), making it hard for short-run players to reward tomorrow actions which are not the commitment action. Second, it ensures short-run players believe the commitment action will be played whenever reputation foundations have bite, which ensures that exit does not “loom large” in the first period (c.f. the discussion succeeding Theorem 1. Markovian beliefs thus rule out (potentially pathological) equilibria where the long-run player is rewarded in the future in a complicated way for revealing they are rational today, even though doing so will cause the game to end with arbitrarily high probability today. The proof is relegated to the PROOF OF PROPOSITION 4 The table below summarizes the main takeaway from the main results in this section and is a core takeaway of the paper.
non-Markov Markov
Reputation Payoff Set Ambiguous (Proposition 4) Stackelberg Payoff (Theorem 1)
No Reputation Folk Theorem (Proposition 2) Static Nash Payoff (Proposition 3)
5 Regime Change
5.1 The Global Games Model
Consider the problem faced by an authoritarian government (player $L$) who must decide how cruelly they should punish dissidents who engage in a revolution. In each period, having successfully defended against an attack, the government can choose to punish dissenters by decreasing their per-period utility by c-c at a cost of $\kappa(c)$, where $\kappa$ is a continuous function satisfying $\kappa(0)=0$. Let $C=[0,\bar{c}]$ and moreover suppose $\kappa(\bar{c})>1$, so that the government never chooses the maximally feasible value of punishment [^13] Note $C=[0,\bar{c}]$ is distinct from the assumption made in the main model that $C$ is countable. This is useful in defining $c^{*}(1,0)$, after which it is sufficient to assume $C=\{0,c^{*}(1,0)\}$ for the remainder of the application. [^14] This is an Inada-type condition that ensures the government’s problem is always well-defined.
The (unit mass, continuum) dissidents now face the following problem. In each period, they know the government is able to fend off $\theta$ attackers, where $\theta\sim P_{\theta}$ is drawn independently in each period, with $\mathbb{E}_{P_{\theta}}[\theta]=\mu$ and $\text{var}(P_{\theta})=\tau^{2}$ [^15] In practice, I will consider the case where $\tau^{2}$ is particularly small. This can be interpreted as a world where the regime has strength $\mu$ and experiences a small idiosyncratic shock in each period to that strength.. Each dissident $i$ cannot see $\theta$ perfectly, instead observing an imperfectly signal $x_{i}$ of $\theta$ with some idiosyncratic noise, $x_{i}=\theta+\varsigma\varepsilon_{i}$, where $\varepsilon_{i}\sim F_{\varepsilon}$ is i.i.d. across each player. Suppose players’ information sets are sufficiently well-behaved that they satisfy the following regularity conditions:
Assumption 3 (Informational Regularity).
The following hold.
  1. (1)
    $F_{\varepsilon}$ and $P_{\theta}$ admit strictly positive, bounded continuous densities; $F_{\varepsilon}$ is symmetric around 0, and $P_{\theta}$ is symmetric around $\mu\in(0,1)$.
  2. (2)
    The first-order conditional distribution $F_{\theta}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
  3. (3)
    The second-order conditional distribution $F_{x}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
Here, $F_{\theta}(\cdot|x)$ is the first order belief over fundamentals of players supposing they see signal $x$ and have prior $P_{\theta}$; $F_{x}(\cdot|x)$ is the second-order conditional belief about other players’ signals. The symmetric assumption on $F_{\varepsilon}$ allows it to be interpreted as white noise. I require $\mu\in(0,1)$ so that the state is not in a dominance region.
Assumption 3 also guarantees $F_{x}$ admits a strictly positive density and that $F_{x}(x|\theta)$ is decreasing in the realized state; moreover, $F_{\theta}(\cdot|x)$ and $F_{x}(\cdot|x)$ are differentiable in the noise term is additive. These smoothness assumptions ensure a first order approach for players is valid, and generalize known conditions standard in the global games literature [^16] I am not aware of a global games setting that does not satisfy these assumptions; for example it nests normal-normal information structures..
Having seen their private signals, the dissidents independently and simultaneously choose in each period an action $a_{i}\in\{0,1\}$, with $a=A=\int a_{i}di$ equal to the total mass of agents taking the aggressive action (i.e. revolt), $a_{i}=1$. If enough dissidents defect (i.e. $A\geq\theta$), then the revolt is successful, the regime is overthrown, and the game ends. The continuation probability $f(A,c)$ is exactly the probability the revolt is successful, and is given by $f(A,c)=\mathbb{P}_{\theta}(A\geq\theta)$ and is independent of $c$.
Player’s payoffs are given by
[…]
Note here payoffs mirror standard global games payoffs (see i.e. Frankel et al. (2003)) where $\theta$ is the peg of the regime, but with the cost of failed coordination (i.e. the coordination friction) flexibly set by the regime at some cost. Standard arguments from the global games literature imply the following.
Proposition 5.
Suppose $\varsigma<\tilde{\varsigma}$ for some $\tilde{\varsigma}>0$ sufficiently small. Then for all $c$, there exists functions $x^{*},\theta^{*}:\mathbb{R}_{+}\to\mathbb{R}$ such that
  1. (1)
    The strategy $s(x,c)=\mathbf{1}\left\{x<x^{*}(c)\right\}$ is the (essentially) unique strategy surviving iterative deletion of strictly dominated strategies.
  2. (2)
    $x^{*}(c)\in\mathcal{C}^{1}$, and is strictly decreasing in $c$.
  3. (3)
    $\theta^{*}(c)\in\mathcal{C}^{1}$, is strictly decreasing in $c$, and $\int s(x,c)dF_{x|\theta}\geq\theta$ if and only if $\theta\leq\theta^{*}(c)$.
Proposition 5 implies not only that $\mathcal{N}(c)$ is single-valued, but that both the set of states which successfully guard against attacks and the cutoff state are themselves monotone around a cutoff, whose evolution is well-behaved. These properties will be useful in the proof of 2 and is proven in the PROOF OF PROPOSITION 5
Property (3) along with the existence of dominance regions implies that Assumption 1 is satisfied. Moreover, since $f(a,c)v(a,c)=\mathbb{P}_{\theta}(A\geq\theta)\cdot(1-\kappa(c))$ and $A^{*}(0)=1$ (i.e. all players attack since it is the weakly dominated strategy), Assumption 2 is satisfied as well.
Finally, the game is clearly monotone when $A$ is ordered in the standard way in $\mathbb{R}$, and satisfies the conditions of Corollary 1. Hence the globalized game satisfies the assumptions of the general model.
5.2 Selecting Equilibria
When is it that the threat of cruelty is enough to deter revolution? Since the game is monotone and satisfies the conditions of Corollary 1, asymptotic $c^{*}$-incentive compatibility is both necessary and sufficient for the regime to deter revolution (as much as possible) in the unique Markovian equilibrium. Rewriting the incentive constraints to the simpler setting of the global game described above yields that asymptotic $c^{*}$-incentive compatibility is equivalent to the inequality
$\delta[V^{*}(\delta)-\underline{V}_{0}(\delta)]>\kappa(c^{*})$
holding for all $\delta$ large and the optimal deterrent choice $c^{*}$. In the global games setting, asymptotic $c^{*}$ incentive compatibility is related to the degree of dispersion of higher order beliefs around the risk-dominant threshold at the average, i.e. whether $\mu>(<)\frac{1}{1+c^{*}}$. Here, recall value $\frac{1}{1+c^{*}}$ is the risk-dominant signal threshold selected by the global game as signals become arbitrarily precise, i.e. $\lim\limits_{\varsigma\to 0}x^{*}(c^{*})=\frac{1}{1+c^{*}}$ by Frankel et al. (2003).
What is the right commitment posture $c^{*}$ to consider when comparing to $\mu$? There are two concerns. First, since $c^{*}(\delta,\varsigma)$ varies with the discount rate and noise precision, the risk-dominant selection threshold varies with the discount rate as well. Second, the underlying probability distributions $(F_{\varepsilon},P_{\theta})$ will also affect the optimal choice for the regime and affect their commitment payoff. To resolve this problem, consider the set
$c^{*}(1,0)=\limsup_{\delta\to 1}\left(\limsup_{\varsigma\to 0}c^{*}(\delta,\varsigma)\right)$
where $\limsup A_{n}$ are the subsequential limits of all selections from $A_{n}$. That $c^{*}(1,0)$ is nonempty follows from two compactness arguments, first by noting $c^{*}(\delta,0)$ is well-defined and nonempty for each $\delta$ and hence so much $c^{*}(1,0)$. Each element of $c^{*}(1,0)$ is approximately optimal for all $\delta\geq\bar{\delta}$ close enough to $1$ and $\varsigma<\bar{\varsigma}(\delta)$, depending potentially on choice of $\delta$, close enough to 0. Throughout, I will fix some $c^{*}\in c(1,0)$ and appeal to continuity of the problem in parameters to make the argument that the same statement holds for $c^{*}(\delta,\varsigma)$ when the first parameter is close to the 1 and the second vanishes. Because the proof of Theorem 1 extends to any arbitrary fixed $c^{*}$, which need not vary with $\delta$ [^17] Theorem 1 is written with flexible commitment postures to compare equilibrium to the commitment payoffs which vary with the discount rate., methods there extend naturally to fixed postures in $c^{*}(1,0)$.
Throughout the statement of Theorem 2, suppose $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ for some $c^{*}\in c^{*}(1,0)$. I make the additional nontriviality [^18] This assumption can be derived by putting mild regularity conditions on $\kappa$, for example that $\kappa^{\prime}(0)=0$. assumption that $c^{*}\neq 0$, so that the commitment type is distinguishable from one playing static Nash.
Theorem 2.
For all $\varepsilon>0$, there exists $\bar{\tau}>0$, $\bar{\delta}<1$, and $\bar{\varsigma}(\tau,\delta)$ vanishing as $\tau\to 0,\delta\to 1$, such that if $\text{var}(P_{\theta})=\tau<\bar{\tau}$, $\delta>\bar{\delta}$, and $\varsigma<\bar{\varsigma}(\tau,\delta)$, then
  1. (1)
    If $R(\mu)<0$, asymptotic $c^{*}$ incentive compatibility holds, $\underline{U}_{0}^{M}(\delta)\geq V^{*}(\delta)-\varepsilon$ and $f(a^{*}(c^{*}))>1-\varepsilon$.
  2. (2)
    If instead $R(\mu)>0$, the $\max\{\underline{U}_{0}(\delta),\bar{U}^{M}(\delta)\}<\varepsilon$.
Theorem 2 highlights a stark difference between the regime’s Stackelberg payoff depending on where their average regime strength when (1) the average becomes an arbitrarily precise summary of the state in each period, (2) the regime becomes arbitrarily patient, and finally (3) the global games perturbation vanishes. Importantly, it delivers a novel implication of risk dominance that is absent from the static game: the regime can only successfully deter rebellion if and only if they are, on average, able to deter rebellion at the risk-dominant level. Should they not be able to do this, then exit immediately happens with high probability.
The result stands in stark contrast to two seminal papers in the literature on dynamic global games with exit. First, it augments the model of Angeletos et al. (2007) to (1) small i.i.d. shocks to the state $\mu$ and (2) reputational incentives for the regime and shows how these perturbations can select a version of the risk-dominant equilibrium (and survival only if $\mu>\frac{1}{1+c^{*}}$) as the unique Markovian equilibrium. Finally it augments the reputational global game model of Huang (2017) to allow for small i.i.d. shocks and a rich exit process, and shows how the regime can fail if their average strength is too weak. This stands in stark contrast to his main result, where there are no attacks that ever occur on the equilibrium path. The differences in these results suggest that taking exit incentives seriously in reputational models can lead to rich results which are more indicative of real-world phenomena, where successful attacks do happen if speculator’s sense the regime is “likely weak.”
The intuition behind the forward direction of Theorem 2 follows from Theorem 1 and a standard concentration inequality argument: if $\mu>\frac{1}{1+c^{*}}$, then the cutoff state as the noise vanishes is below the mean, so as the variance vanishes the regime must survive most of the time. Conversely, if $\mu<\frac{1}{1+c^{*}}$, then most of the time the regime dies, and hence they must attain a low payoff. A formal proof can be found in the PROOF OF THEOREM 2
6 Discussion
This paper studied a model of reputation where players’ actions endogenously affected the probability of exit in the game (and hence the effective discount factor was endogenous). Such endogeneity creates several difficulties for the analysis of the repeated game. First, standard ways of normalizing payoffs no longer apply, and thus players’ values from different actions can vary directly with the discount rate. Second, folk theorems need no longer hold because the future can be “small” relative to actions today, making enforceability infeasible. Third, the addition of reputation effects need not select the long-run player’s Stackelberg payoff even when the long-run player becomes arbitrarily patient
I identify two conditions that help ameliorate these problems. These conditions are an incentive compatibility condition on the effective rate of future discounting: the long-run player wishes to play the Stackelberg action if recommended to do so, supposing maximal variation in continuation payoffs after; and a credible belief condition, represented by the Markovian refinement: short-run players believe the long-run player has an incentive to build a reputation when those incentives are present. Applied to a global games setting, these conditions yield new predictions about the survival of risk dominant equilibria in a dynamic regime change game.
There are several fruitful directions for future analysis. First, understanding the limits of endogenous exit when players’ actions are imperfectly monitored is a natural next step. The methods developed in this paper need not apply because reputation formation and incentive provision are much more subtle with imperfect monitoring, and the standard law-of-large-numbers intuitions do not apply with endogenous exit. Second, understanding the case with multiple commitment types is also promising, away from sequential equilibria—here, the set of payoffs may be pathological because of continuation payoffs at off-path deviations (i.e. taking a committed action that is not on the path of play). Understanding the set of Nash equilibria in this game and their payoff bounds is potentially promising. Finally, there is scope to apply the machinery in this paper to several other settings. For example, in the infinite chain store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982), my findings can give predictions about when the long-run player can deter entry even when entry affects the probability the incumbent firm must declare bankruptcy. Alternatively, Theorem 1 can be used to give conditions under which Markovian equilibria have reputation foundations in stopping time games where short-run or myopic players can stop at any time (i.e. adapted versions of Chassang (2010)). Finally, outside of Markovian predictions, these results may give new interpretations of “bad reputation” style results in the spirit of Ely and Valimaki (2003) and Ely et al. (2008).
Appendix: Omitted Proofs
Proof of Lemma 1
Proof.
Let $\{\bar{f}_{\tilde{t}}\}\subset\{f_{t}\}$ be a subsequence such that $\lim\limits_{\tilde{t}\to\infty}\bar{f}_{\tilde{t}}=\bar{f}<1-\eta$, for some $\eta>0$ fixed; moreover, by truncating finitely many terms at most, suppose $\bar{f}_{\tilde{t}}<1-\eta$ for all $s\in\mathbb{N}$. Now for any $T\in\mathbb{N}$, define the number of times that $f_{t}<1-\eta$ to be $s(T)=\#\{s:f_{s}<1-\eta\}$. By assumption of $\eta$, we know $\lim\limits_{T\to\infty}s(T)=\infty$. We now have that $\prod_{s=0}^{\infty}f_{t}<(1-\eta)^{s(T)}$. Letting $\bar{v}$ be chosen so that $|v_{t}|\leq\bar{v}$ for all $t$, we have
$0\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}\delta^{t}v_{t}\prod_{s=0}^{\infty}f_{s}\right|\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}\delta^{t}\bar{v}(1-\eta)^{s(t)}\right|$
$\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{T}\delta^{t}\bar{v}(1-\eta)^{s(t)}\right|+\left|(1-\delta)(1-\eta)^{s(T)}\sum_{t=T+1}^{\infty}\delta^{t}\bar{v}\right|$
$\leq(1-\eta)^{s(T)}\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=T+1}^{\infty}\delta^{t}\bar{v}\leq(1-\eta)^{s(T)}\bar{v}.$
Since this holds for all $T\in\mathbb{N}$, taking $T\to\infty$ (and hence $s(T)\to\infty$) gives the argument. ∎
Proof of Lemma 2
Proof.
Note
$\sum_{t=0}^{\infty}\delta_{1}^{t}b_{t}=\delta^{T}\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}\geq 0\iff\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}$
As a function of $\delta$, this series is absolutely convergent for all $\delta\in(0,1)$ and the partial sums converge uniformly on this interval so we can differentiate under the summation: thus,
$\frac{d}{d\delta}\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}=\sum_{t=0}^{\infty}(t-T)\delta_{1}^{t-T-1}b_{t}<\infty$
where we use the fact $b_{t}$ is bounded to guarantee the sum is well-defined. Note each term in this summation is nonnegative. If $t<T$, then tT<0t-T<0 and $b_{t}<0$, so $(t-T)\delta^{t-T-1}b_{t}>0$, while if $t>T$, then $(t-T)\delta^{t-T-1}\geq 0$ and $b_{t}\geq 0$ as well. Moreover, since $\{b_{t}\}$ is not identically zero, at least one term in the sum above must be strictly positive, and hence the derivative is strictly positive. We thus have that
$\sum_{t=0}^{\infty}\delta_{2}^{t-T}b_{t}>0\implies\delta_{2}^{T}\sum_{t=0}^{\infty}\delta_{2}^{t-T}b_{t}=\sum_{t=0}^{\infty}\delta_{2}^{t}b_{t}>0$
where we use the fact $\delta_{2}>\delta_{1}$, and finally multiply by $\delta_{2}^{T}>0$. ∎
Proof of Proposition 2
Proof.
Some preliminaries. First, the limit exists since as $\delta$ increases the payoff set weakly expands. Moreover, $\underline{V}_{0}(\delta)$ must be the long-run player’s lowest possible equilibrium payoff as 0 is dominant in the stage game. Similarly, $V^{*}(\delta)$ must be their highest equilibrium payoff because short-run players must myopically best reply. Next, since $\mu_{0}(h^{t})=1$, we consider assessments where $\mu(h^{t})(\omega_{R})=1$ always and find subgame perfect equilibrium. Finally, throughout the proof, assume players have access to a public randomization device so that the payoff set is convex.
It is clear that $\underline{V}_{0}(\delta)$ is attainable by an equilibrium where the lowest stage-game Nash equilibrium is played at every history. Suppose now the second condition is satisfied. Consider the equilibrium with the following strategies.
  1. (1)
    If $h^{t}=\{c^{*}\}_{s=0}^{t-1}$, $\sigma_{L}(h^{t})=c^{*}$, $\{\sigma_{i}(h^{t})\}_{i\in[N]}=a^{*}\in\mathcal{N}(c^{*})$.
  2. (2)
    At any other history, $\sigma_{L}^{*}(h^{t})=0$ and $\{\sigma_{i}(h^{t})\}_{i\in[N]}=\underline{a}\in\mathcal{N}(0)$, with $\underline{a}$ chosen to attain the lowest Nash equilibrium payoff $\underline{V}_{0}(\delta)$.
By construction short-run players are myopically best replying. Clearly at histories where $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$ the long-run player’s strategy is sequentially rational. At a history where only $c^{*}$ has been played, Assumption 2 implies the long-run player’s most profitable deviation is to play 0; this is unprofitable so long as
$f(a^{*},c^{*})[v(a^{*},c^{*})+\delta V^{*}(\delta)]\geq f(a^{*},0)[v(a^{*},0)+\delta\underline{V}_{0}(\delta)].$
Divide through by $f(a^{*},0)$ and collect terms in $\delta$ on the left hand side to conclude that the above expression is equivalent to
$\delta[L(a^{*},c^{*})V^{*}(\delta)-\underline{V}_{0}(\delta)]\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*})$
as desired. Thus, so long as $\delta>\bar{\delta}$, $\underline{U}(\delta),\bar{U}(\delta)=[\underline{V}_{0}(\delta),V^{*}(\delta)]$.
The converse. Suppose the second condition fails. Some algebra implies it must be that for any possible continuation payoffs $V(c^{*}),V(0)$ for playing $c^{*}$ and 0 respectively, we have
$f(a^{*},c^{*})[v(a^{*},c^{*})+\delta V(c^{*})]<f(a^{*},0)[v(a^{*},0)+\delta V(0)]$
and so it is never possible to induce the long-run player to play $c^{*}$ on path. Thus the long-run player’s largest payoff is bounded away from $\underline{V}_{0}^{*}(\delta)$ for any $\delta$. ∎
Proof of Proposition 3
Proof.
First, note that if $\mu_{0}(\omega_{R})=1$, then it must be that in every sequential equilibrium $\mu(h^{t})(\omega_{R})=1$ as well (as the agent has a degenerate prior belief). Now fix any history $h^{t}$; at any successor history $h^{s},h^{s^{\prime}}\succ h^{t}$, it must be that $\sigma(h^{s})=\sigma(h^{s^{\prime}})$, so in particular future play cannot vary based on play today. Thus, the long-run player’s action must be myopically optimal, and hence $\sigma_{L}(h^{t})=0$ at any history $h^{s}$. Short-run players must play the same myopic best response at each history, so in equilibrium $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(0)$ always. Hence all equilibria are repeated static Nash. Public randomization convexifies the payoff set. ∎
Proof of Lemma 3
Proof.
Let $(\sigma,\mu)$ be a sequential equilibrium and let $(\sigma^{n},\mu^{n})$ be the fully supported strategy which converges to it. Now fix any history $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$ noting it is supported by $\mathbb{P}^{\sigma^{n}}$ for all $n$. It must be that $\mu^{n}(h^{t})(\omega_{R})=1$, since no committed type could have produced this history but there is positive probability the rational type did. This implies $\lim\limits_{n\to\infty}\mu^{n}(h^{t})(\omega_{R})=\mu(h^{t})(\omega_{R})=1$, finishing the proof. ∎
Proof of Lemma 4
Proof.
Let $(\sigma,\mu)$ be Markov perfect. By Proposition 3, any continuation strategy at histories $h^{t}$ where $\mu(h^{t})(\omega_{R})$ must be a repeated static Nash profile, so in particular $\sigma_{L}(h^{t})=0$.
Consider now a history $h^{t}$ where $\mu(h^{t})(\omega_{R})<1$ (noting these are all on path), and suppose the lemma is false. Then there exists some on-path continuation history $h^{t}$ where at $h^{t}\succ h^{t-1}$, $h^{t-1}=\{c^{*}\}_{t=0}^{t-2}$ but $h_{t}=c\not\in\{0,c^{*}\}$. Thus, $\mu(h^{t})(\omega_{R})=1$ and the long-run player’s continuation value at $h^{t}$ must be the same as their continuation value at $\{h^{t-1},0\}$. But then playing 0 instead of $c$ for sure is a strictly profitable deviation by Assumption 2, contradicting sequential rationality of $(\sigma,\mu)$. ∎
Proof of Lemma 5
Proof.
Suppose not. Let $(\sigma,\mu)$ be an MPE and let $h^{t}$ be a history at which the long-run player mixes between 0 and $c^{*}$. By the argument in Lemmas 3 and 4, this can only possibly occur on a history $h^{t}=\{c^{*}\}_{s=0}^{t-1}$ for some $t$; call such a history the time-$t$ clean history. Define
$\mathcal{T}=\{t:h^{t}\text{ is clean and }\sigma_{L}(h^{t})\text{ mixes.}\}$
First, $|\mathcal{T}|$ cannot be infinite. Suppose not, so there is a sequence of times $\{t_{n}\}_{n=1}^{\infty}\in\mathcal{T}$. Note that when the long-run player mixes, both (1) $h^{t_{n}}\in\mathbb{P}^{\sigma}$ where $h^{t_{n}}$ is clean, and (2) the long-run player’s action is statistically identifiable compared to the committed type. This implies by a concentration of measure argument (see Mailath and Samuelson (2006) Chapter 15) that $\lim\limits_{n\to\infty}\mu(h^{t_{n}})(\omega_{c^{*}})=1$. Because short-run player actions are upper hemi-continuous in $\mu(h^{t_{n}})$, there must exist $N$ sufficiently large that for all $n\geq N$, $\mathcal{V}(h^{t_{n}}|\sigma)\geq\mathcal{V}(c^{*},h^{t_{n}})\geq V_{*}(\delta)-\varepsilon$ for some $\varepsilon>0$, where $\mathcal{V}(c^{*},h^{t})$ is the long-run player’s continuation history from simply always taking $c$ in the future and $\mathcal{V}(h^{t}|\sigma)$ is their continuation value from $\sigma$. Yet mixing implies the long-run player must be indifferent between taking 0 (and obtaining at most $\bar{V}_{0}(\delta)$ hereafter) and taking $c^{*}$ (and obtaining that future continuation value). In particular, take $N$ sufficiently large and $a=\{\sigma_{i}(h^{t_{N}})\}$ that
$f(a,0)[v(a,0)+\delta\mathcal{V}(\{h^{t},0\}|\sigma)]=f(a,c^{*})[v(a,c^{*})+\delta V(\{h^{t_{n}},c^{*}\}|\sigma)].$
Yet we also know from Proposition 3 that
$f(a,0)[v(a,0)+\delta\mathcal{V}(\{h^{t},0\}|\sigma)]\leq f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]$
and from our observation above that
$f(a,c^{*})[v(a,c^{*})+\delta V(\{h^{t_{n}},c^{*}\}|\sigma)]\leq f(a,c^{*})[v(a,c^{*})+\delta(V_{*}(\delta)-\varepsilon)].$
Yet the assumption the long-run player asymptotically satisfies $c^{*}$ IC implies
$f(a,0)[v(a,0)+\delta\bar{V}(\delta)]<f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
for $a$ since $a\in\mathcal{N}(\alpha c)$. Taking everything together and using asymptotic $c^{*}$-incentive compatibility implies for all $N$ sufficiently large that
$f(a,0)[v(a,0)+\delta\mathcal{V}(\{h^{t},0\}|\sigma)]<f(a,c^{*})[v(a,c^{*})+\delta V(\{h^{t_{n}},c^{*}\}|\sigma)]$
contradicting the indifference condition which is necessary to mix.
Thus, $|\mathcal{T}|$ is finite. Let $T>0$ be the last time at which the long-run player mixes at a clean history. This implies after $T$, the long-run player attains at least $V_{*}(\delta)$ when they play 0. Thus, for $a=\{\sigma_{i}(h^{T})\}\in\mathcal{N}(\alpha c)$ for some $\alpha\in[0,1]$, we have
$f(a,0)[v(a,0)+\delta\underline{V}_{0}(\delta)]=f(a,c^{*})[v(a,c^{*})+\delta V^{*}(\delta)]$
This again contradicts asymptotic $c^{*}$ IC. Hence the long-run player never mixes. ∎
Proof of Theorem 1 (Converse)
Proof.
Suppose the condition fails and fix any $\delta\in(0,1)$. There are three cases. First, suppose the inequality fails strictly at every $\alpha\in[0,1]$. Then in particular
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]>f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
This incentive compatibility constraint implies that the rational long-run player would always prefer to separate and play 0 (starting from any time-0 belief) if they are guaranteed $\bar{V}_{0}(\delta)$ afterwards. In particular, it implies the following is a Markovian equilibrium:
  1. (1)
    The rational long-run player plays 0 in every period.
  2. (2)
    short-run players play $a_{*}=a^{*}(c^{*})$ at all $h^{t}$ where $\mu(h^{t})(\omega_{R})=0$, with $a_{*}$ chosen to deliver the long-run player expected continuation payoff $V_{*}(\delta)$.
  3. (3)
    If $\mu(h^{t})(\omega_{R})=1$, play $a_{0}^{*}\in\mathcal{N}(0)$ chosen to deliver the long-run player an expected continuation payoff $\bar{V}_{0}(\delta)$.
  4. (4)
    At histories where $\mu(h^{t})(\omega_{R})=\mu_{0}(\omega_{R})$, play any $a\in\mathcal{N}(\mu_{0}(\omega_{c^{*}})c^{*})$.
Second, suppose instead that the inequality fails for some $\alpha$ but that it holds for $\alpha=1$. This implies first that it is possible to sustain the Markovian equilibrium where the long-run player plays $c^{*}$ at all histories. Moreover, there must exist some $\bar{\alpha}\in(0,1)$ with $a\in\mathcal{N}(\alpha c^{*})$ such that
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]=f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
by the intermediate value theorem, since the incentive compatibility constraint is continuous in $\alpha$. This implies that there exists a Markovian equilibrium of the following form:
  1. (1)
    If $\mu(h^{t})(\omega_{R})\neq\{\mu_{0}(\omega_{R}),1\}$, then have $\sigma_{L}(h^{t})=c^{*}$ and $\{\sigma_{i}(h^{t})\}\in\mathcal{N}(c^{*})$ be chosen to give the long-run player $V_{*}(\delta)$.
  2. (2)
    If $\mu(h^{t})(\omega_{R})$, play static Nash $\sigma(h^{t})$ giving the long-run player $\bar{V}_{0}(\delta)$.
  3. (3)
    If $\mu(h^{t})(\omega_{R})=\mu_{0}(\omega_{R})$, have $a=\{\sigma_{i}\}\in\mathcal{N}(\alpha c*)$ chosen so that
    $f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]=f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]$
    and $\sigma_{L}(h^{t})\in(0,1)$ chosen so that
    $\mu_{0}(\omega_{R})\sigma(h^{t})(c^{*})+(1-\mu_{0}(\omega_{R}))=\alpha$
    Noting this is well-defined so long as $\mu_{0}(\omega_{R})>1-\alpha$.
It is clear that at all histories, the short-run player is best replying, as is the long-run player at all histories outside of $\varnothing$. Moreover, at the first history $h^{t}=\varnothing$, by choice of $\alpha$, the long-run player is indifferent between $c^{*}$ and 0 and thus is willing to randomize given $a$. Hence this is a Markov perfect equilibrium. This gives the long-run player an equilibrium payoff of
$f(a,0)[v(a,0)+\delta\bar{V}_{0}(\delta)]<f(a_{*},0)[v(a_{*},0)[v(a_{*},0)+\delta\bar{V}_{0}(\delta)]]$
$\leq f(a_{*},c^{*})[v(a_{*},c^{*})+\delta V_{*}(\delta)]$
where the first inequality follows from monotonicity of the game and the second follows from the fact the incentive compatibility condition holds at $\alpha=1$ where $a_{*}=a^{*}(c^{*})$ is the largest element of $N(c^{*})$, since by the monotonicity assumption this is the one which minimizes the stage-game payoff. Taking $\delta\to 1$ implies that the lowest equilibrium payoff is still separated from the commitment payoff by at least $f(a_{*},0)v(a_{*},0)-f(a,0)v(a,0)>0$; since this gap is independent of $\delta$, it implies the result.
Finally, suppose the inequality fails when $\alpha=1$. Then there cannot be an equilibrium where the long-run player imitates $c^{*}$ for sure. The only other pure Markovian equilibrium gives the long-run player their minimax payoff and implies the result. If instead the long-run player mixes in a Markovian equilibrium at some on-path history, they secure (for some $a\in\mathcal{N}(\alpha c^{*})$ and some $\alpha\in(0,1)$) a payoff of at most
$f(a,c^{*})[v(a,c^{*})+\delta V_{*}(\delta)]<f(a^{*},c^{*})[v(a^{*},c^{*})+\delta V_{*}(\delta)]=V_{*}(\delta)$
by monotonicity and hence must secure a payoff smaller than $V_{*}(\delta)$. Taking $\delta\to 1$ finishes this last case and thus the proof. ∎
Proof of Corollary 1
Proof.
First, note that under these conditions, the incentive compatibility constraint in the folk theorem can be written as
$\delta[V^{*}(\delta)L(c^{*})-\underline{V}_{0}(\delta)]\geq v(0)-v(c^{*})L(c^{*})$
and asymptotic $c^{*}$ incentive compatibility can be written as
$\delta[V^{*}(\delta)L(c^{*})-\underline{V}_{0}(\delta)]>v(0)-v(c^{*})L(c^{*})$
because dependence on $a$ is suppressed in the likelihood ratio and $|\mathcal{N}(0)|$ is a singleton so $\underline{V}_{0}(\delta)=\bar{V}_{0}(\delta)$. These are the same incentive compatibility constraints modulo the equality constraint. But because in the folk theorem the right hand side is constant in $\delta$ but the left hand side is increasing, there must exist $\delta$ large enough where the first constraint holds strictly. Hence the two statements are equivalent.
That there is a unique Markovian equilibrium when the IC constraint is satisfied then follows by noting that (1) there is a unique static Nash equilibrium (and hence a unique strategy at all histories in an MPE where $\mu(h^{t})(\omega_{R})=1$) and (2) that $\Omega(\delta)=\{0,c^{*}(\delta)\}$ contains only two elements, so any histories where $\mu(h^{t})(\omega_{R})<1$ must be clean with a fixed element. Thus the unique MPE is the one that plays $c^{*}(\delta)\in\Omega(\delta)$ at every history. ∎
PROOF OF PROPOSITION 4
Proof.
Fix $\varepsilon>0$ and preferences $(u,v)$. Fix some action $c$ and $a^{*}(c)$ be its lower Stackelberg best response where $v(a^{*}(c),c)>0$, which exists by the nontriviality assumption. Set $f(a^{*}(c),c)=1-\eta,f(a^{*}(c),0)=\eta$. Let $\underline{a}=a^{*}(0)\in\mathcal{N}(0)$ be such that $(\underline{a},0)$ gives the long-run player payoff $\underline{V}_{0}(\delta)$. Set moreover $f^{*}(\underline{a},0)=f^{*}(\underline{a},c)=\eta$. Note for $\eta$ sufficiently small, playing $c$ must be preferred to playing 0 since
$\frac{v(a^{*}(c),c)(1-\eta)}{1-\delta(1-\eta)}\geq\frac{v(a,0)\eta}{1-\delta\eta}$
for any $a\in\mathcal{N}(0)$. This in particular implies for all $\eta>0$, there is an optimal commitment posture $c^{*}\neq 0$ and in particular one which guarantees a payoff of at least $v(a^{*}(c),c)>0$ in each period conditional on surviving.
Let $c^{*}(\delta,\eta)$ be the optimal commitment posture given discount rate $\delta$ and exit function given as above (and $V_{*}(\delta,\eta)$ its associated Stackelberg payoff). In this case, Proposition 2 implies the folk theorem is equivalent to
$f(a^{*}(c^{*}(\delta,\eta)),c^{*})[v(a^{*}(c^{*}(\delta,\eta)),c^{*})+\delta V_{*}(\delta,\eta)]\geq f(a^{*},0)[v(a^{*}(0)+\delta\underline{V}_{0}(\delta)]$
The term inside the brackets on the right hand side is independent of $\eta$, and the left hand side by assumption is bounded from below by
$f(a^{*}(c),c)[v(a^{*}(c),c)+\delta V_{*}(\delta,\eta)]$
as $\eta$ vanishes. But then our condition is implied by the requirement that
$\frac{f(a^{*}(c),c)}{f(a^{*},0)}[v(a^{*}(c),c)+\delta V_{*}(\delta,\eta)]\geq[v(a^{*}(0)+\delta V_{*}(\delta)]$
$\iff$ $\frac{1-\eta}{\eta}[v(a^{*}(c),c)+\delta V_{*}(\delta,\eta)]\geq[v(a^{*}(0)+\delta\underline{V}_{0}(\delta)]$
which is clearly true as $\eta\to 0$. Hence under this technology asymptotic $c^{*}$ optimality is satisfied for all $\eta$ sufficiently small.
Fix now some $\eta$ sufficiently small for the folk theorem to hold (in particular so that $V^{*}(\delta)$ can be sustained as an equilibrium) under $f$ and consider the following (non-Markovian) strategy profile at time $t=0$:
  1. (1)
    Have the long-run player play 0 in the first period.
  2. (2)
    Have the short-run player’s play to $a^{*}(\mu_{0}(\omega_{c^{*}}c^{*}))$, i.e. best respond to their prior belief supposing the rational type plays 0.
  3. (3)
    Regardless of the long-run player’s action, given them a payoff of $V^{*}(\delta)$ tomorrow.
Clearly $V^{*}(\delta)$ is sustainable by the folk theorem tomorrow; and because the rational long-run player’s action today does not affect their future continuation payoff, they will play their dominant action 0. Finally, Condition (2) is exactly the condition for short-run players to myopically best reply to this strategy by the long-run player. As $\mu_{0}(\omega_{R})\to 1$, short-run players (by upper hemi-continuity) converge to an action in $\mathcal{N}(0)$. In particular, this implies $\underline{a}$ must be short-run players a (weakly) lower payoff than $\lim\limits_{\alpha\to 0}v(a^{*}(\alpha c^{*}),c^{*})$.
This equilibrium strategy profile gives the long-run player a payoff of exactly
$f(\underline{a},c)[v(\underline{a},0)+\delta V_{*}(\delta)]=\eta[v(\underline{a},0)+\delta V_{*}(\delta)].$
Choosing $\eta<\frac{\varepsilon}{v(\underline{a},0)+\delta V_{*}(\delta)}$ then implies the argument. ∎
PROOF OF PROPOSITION 5
Proof.
Fix a value $c$, and note the induced stage-game is a finite monotone (decreasing) supermodular game with dominance regions $\infty,0)$ and $(1,\infty]$. This implies by the proof of Theorem 1 of Frankel et al. (2003) that for all $\varsigma$ sufficiently small, $s(x,c)=\mathbf{1}\left\{x<x^{*}(c)\right\}$ is the unique equilibrium strategy profile.
Next, let $U_{i}(c,x,x^{*})$ be the interim utility differential between taking $a=1$ and $a=0$ for an agent observing signal $x$ when players use a cutoff strategy $x^{*}$ and the cost of failed revolution is $c$. In equilibrium, it must be that $U_{i}(c,x^{*}(c),x^{*}(c))=0$. To ensure that $x^{*}(c)$ is differentiable and strictly increasing, we appeal to the implicit function theorem; some algebra gives that $x^{*}(c)$ is implicitly defined as the solution to the equation
$U_{i}(c,x,x)=(1+c)F_{\theta}(F_{x}(x|x)|x)-c=0.$
We have that
$\frac{\partial U_{i}}{\partial c}=F_{\theta}(F_{x}(x|x)|x)-1\text{ and }\frac{\partial U_{i}}{\partial x}=(1+c)\left(F_{\theta}(F_{x}(x|x)|x)\cdot f_{x}(x|x)\cdot\frac{\partial}{\partial 2}F_{\theta}(F_{x}(x|x)|x)\cdot\frac{\partial}{\partial 2}F_{x}(x|x)\right)$
where $\frac{\partial}{\partial 2}$ refers to the derivative in the conditioning argument. Note the derivative in $c$ is strictly negative as $P_{\theta}(\cdot)<1$ always; meanwhile, the second term is strictly negative by Assumption 3. Thus, the implicit function theorem implies
$\frac{\partial x^{*}(c)}{\partial c}=-\frac{\partial U_{i}}{\partial c}\left(\frac{\partial U_{i}}{\partial x}\right)^{-1}<0,$
as desired. The last statement follows similarly. Regime change only occurs if $F_{x}(x^{*}(c)|\theta)>\theta$, so $\theta^{*}(c)$ is implicitly defined by the equation $F_{x}(x^{*}(c)|\theta)-\theta=0$. Differentiating yields
$\frac{d}{d\theta}(F_{x}(x^{*}(c)|\theta)-\theta)=\frac{\partial}{\partial 2}F_{x}(x^{*}(c)|\theta)-1<0$
as $\frac{\partial}{\partial 2}F_{x}(x|\theta)<0$ for all $x$; this then implies $\theta^{*}(c)$ is globally $\mathcal{C}^{1}$. Applying the implicit function theorem again implies its derivatives is given by
$\frac{\partial\theta^{*}(c)}{\partial c}=-\left(f_{x}(x^{*}(c)|\theta)\frac{\partial x^{*}(c)}{\partial c}\right)\left(\frac{\partial}{\partial 2}F_{x}(x^{*}(c)|\theta)-1\right)^{-1}<0,$
and so higher costs $c$ decrease the region of states where attacks are successful. ∎
PROOF OF THEOREM 2
For all $\varepsilon>0$, there exists $\bar{\tau}>0$, $\bar{\delta}<1$, and $\bar{\varsigma}(\tau,\delta)$ vanishing as $\tau\to 0,\delta\to 1$, such that if $\text{var}(P_{\theta})=\tau<\bar{\tau}$, $\delta>\bar{\delta}$, and $\varsigma<\bar{\varsigma}(\tau,\delta)$, then
  1. (1)
    If $R(\mu)<0$, asymptotic $c^{*}$ incentive compatibility holds, $\underline{U}_{0}^{M}(\delta)\geq V^{*}(\delta)-\varepsilon$ and $f(a^{*}(c^{*}))>1-\varepsilon$.
  2. (2)
    If instead $R(\mu)>0$, the $\max\{\underline{U}_{0}(\delta),\bar{U}^{M}(\delta)\}<\varepsilon$.
Proof.
Fix some distribution $P_{\theta}$. Note
$\lim\limits_{\varsigma\to 0}F_{x}\left(\frac{1}{1+c^{*}}\bigg{|}x\right)\to\begin{cases}0\text{ if }x<\frac{1}{1+c^{*}}\\ 1\text{ if }x>\frac{1}{1+c^{*}}\end{cases}$
because $F_{x}(\cdot|\theta)$ converges to $\delta_{\theta}$ as $\varsigma\to 0$ (people become arbitrarily sure about the state). Since $\theta^{*}(c)$ is continuous and zeros $F_{x}\left(\frac{1}{1+c^{*}}|\theta\right)-\theta$, this implies $\theta^{*}(c)=\frac{1}{1+c^{*}}$.
The first statement. Suppose $\frac{1}{1+c^{*}}<\mu$ and let $\eta=\mu-\frac{1}{1+c^{*}}$, noting this gap is independent of $\tau$. We want to prove
$\delta(V^{*}(\delta)-\underline{V}_{0}(\delta))>\kappa(c^{*})$
in the right order of limits. Some algebra implies where $\bar{P}_{\theta}(x)=1-P_{\theta}(\theta^{*}(x))$ is the probability of survival given a commitment posture. Taking $\delta$ sufficiently large and noting the inequality is strict implies that this statement is equivalent to the requirement
$\bar{P}_{\theta}(c^{*})-\bar{P}_{\theta}(0)\geq\kappa(c^{*})[1-\bar{P}_{\theta}(0)].$
Note from here that
$\bar{P}_{\theta}(c^{*})=1-P_{\theta}(\theta^{*}(c^{*}))>1-P_{\theta}(\mu-\eta)\geq 1-\frac{\text{var}(P_{\theta})}{\eta^{2}}>1-\frac{\bar{\tau}}{\eta}.$
The middle inequality uses Chebychev’s inequality. Thus $\bar{P}_{\theta}(c^{*})\to 1$ as $\bar{\tau}\to 0$; similarly $\bar{P}_{\theta}(0)\to 0$ as $\bar{\tau}\to 0$; together this implies the incentive constraint holds. Theorem 1 and Corollary 1 imply the regime can sustain uniquely $c^{*}$ and some best response to $c^{*}$ in the unique Markovian equilibrium, as desired. Finally because $\bar{P}_{\theta}(c^{*})=f(c^{*})$, the probability of exit vanishes as $\bar{\tau}\to 1$. This implies both parts of the first statement, noting $\bar{\varsigma}$ can depend on the choice of $(\tau,\delta)$.
Now consider the case where $\frac{1}{1+c^{*}}>\mu$, now setting $\eta=\frac{1}{1+c^{*}}-\mu>0$. Note a necessary condition for incentive constraints to hold is for
$\frac{\bar{P}_{\theta}(c^{*})-\bar{P}_{\theta}(0)}{1-\delta\bar{P}_{\theta}(0)}>\kappa(c^{*})$
For any fixed $\delta$, note that as $\tau$ vanishes, one has that
$1-P_{\theta}(\theta^{*}(\sigma^{*}(h^{t}))\leq 1-P_{\theta}(\theta^{*}(c^{*}))<1-P_{\theta}(\mu+\eta)\leq\frac{\text{var}(P_{\theta})}{\eta^{2}}<\frac{\bar{\tau}}{\eta^{2}}$
again using Chebychev’s inequality. Since $\theta^{*}(c)<\theta^{*}(0)$, we can conclude the numerator must vanish while the denominator converges to $1-\delta$. Hence this inequality will fail strictly for all sufficiently small $\tau$ for any fixed $\delta$. This implies that $c^{*}$-incentive compatibility must fail strictly, so by the converse Theorem 1 in the unique Markovian equilibrium the regime plays 0 in every period. Since $\bar{P}_{\theta}(0)$ vanishes, $\underline{U}_{0}(\delta)<\varepsilon$. Since this was the unique Markovian equilibrium, $\underline{U}^{M}(\delta)<\varepsilon$ as well. This finishes both directions of the proof. ∎