### v5

# Reputation in the Shadow of Exit

Abstract
I study reputation formation in repeated games where player actions endogenously determine the probability the game permanently ends. Permanent exit can render reputation useless even to a patient long-lived player whose actions are perfectly monitored, in stark contrast to canonical commitment payoff theorems. However, I identify tight conditions for the long-run player to attain their Stackelberg payoff in the unique Markov equilibrium. Along the way, I highlight the role of Markov strategies in pinning down the value of reputation formation. I apply my results to give qualified commitment foundations for the infinite chain-store game. I also analyze repeated global games with exit, and obtain new predictions about regime survival.
Keywords: Repeated Games, Reputation, Commitment Payoff Theorem, Endogenous Exit.
JEL Codes: C73, D74, I14.
1 Introduction
Reputational incentives have long been recognized as a powerful way to refine predictions in infinitely repeated games. By endowing the long-run player with the ability to imitate a commitment type who always takes some action $a$, a rational long-run player can build a reputation for playing $a$ and thus secure the payoff they would obtain as-if they were known to be committed to that action. Successful reputation formation relies on two important features of the repeated game: First, that the game itself is repeated sufficiently often, so that the long-run player can undertake the action enough times to build their reputation, and second, that they are extremely patient, so that the “reputation-building” phase of the game contributes little to their average payoff.
Yet there are a variety of settings where the first assumption may be unrealistic. Consider, for example, the chain store game of Kreps and Wilson (1982) and Milgrom and Roberts (1982) (hereafter, simply “chain store game” or “KWMR”). A crucial assumption in their setup is that the incumbent is able to repeatedly maintain their position in an a-priori fixed number of markets, even if they unsuccessfully deter entry. Yet in many incumbent-entrant games, successful entry by a challenger can displace the incumbent, changing the structure of the game in future periods, and either forcing the incumbent to exit or the game to prematurely end [^2] For example, the monopolist loses their market power in the KWMR chain store game if a challenger successfully enters and is accommodated.. Similarly, in the product choice game of Mailath and Samuelson (2006) (Chapter 15.1) and Pei (2020), it may be that a long enough string of low quality forces the firm to exit. Finally, in models of repeated regime change (c.f. Huang (2017)), a successful currency attack could end the current regime, i.e. lead to the toppling of the central bank. Players, realizing this, may change their behavior, which could in turn affect reputational dynamics.
In such games, the assumption the stage game will be infinitely repeated can be unreasonable, as player actions may lead to premature permanent exit. Consequently, the expected number of periods play persists may be much shorter than the “upper bound on the number of times the short-run players can fail to play as Stackelberg followers,” (c.f. Fudenberg and Levine (1989)) blocking both the long-run player’s ability to build a reputation and for the early stages of the game to be unimportant for their average payoff. Indeed, without strong assumptions on the exit technology itself, little can be said about reputation formation in the shadow of exit, restoring the indeterminancy of the folk theorem present in the complete information analogue (see Proposition 3). Despite the potential appropriateness of directly modeling exit in several of the leading examples studied by the reputation literature, little has been said about the effect exit has on incentives for reputation formation.
This paper studies reputation formation in repeated games where player actions endogenously determine if the game permanently ends. Formally, I consider a model where many short-run players (choosing $a$) interact with a long-run player (choosing $c$) who can be either a rational type or a single committed type who plays a Stackelberg action. In addition to payoffs, the action profile $(a,c)$ also jointly determines a survival probability the game permanently ends. The rational type wishes to maximize their expected payoff by potentially imitating the committed type. Crucially, fear of exit can shut down traditional incentives for reputation formation, as the effective discount rate given short-run player actions can be bounded away from $1$ even when the long-run player is very patient. Thus there is no guarantee that investing in the Stackelberg action today will be rewarded with short-run Stackelberg best replies with high frequency in the future, rendering standard arguments from the reputation literature inapplicable.
The main result of this paper identifies two jointly necessary and sufficient conditions that together are enough for reputation effects to select the long-run player’s Stackelberg payoff as the unique Markov equilibrium payoff:
  1. (1)
    Incentive compatibility: the “effective” discount fact is high enough that playing the Stackelberg action can be enforced in some equilibrium of the repeated game.
  2. (2)
    Credible beliefs (a Markov refinement): whenever short-run players hold identical beliefs about the long-run player’s rationality, they act the same way.
The incentive compatibility condition is captured by a single enforceability constraint, which is taken as hypothesis for the statement of Theorem 1 (see also Proposition 1). The credible belief condition is captured by a Markov refinement that short-run player actions are measurable with respect to their belief about the long-run players’ type. This ensures that short-run players act on their beliefs about the long-run player’s rationality, and rules out equilibria where the long-run player exits with high probability in the first period due to a lack of variation in continuation play in the future. Indeed, I prove in the absence of this condition an anti-commitment payoff theorem (Proposition 3) by constructing revelation equilibria exactly of this form which guarantee the long-run player an arbitrarily low payoff absent the Markov refinement, even when payoffs satisfy my incentive compatibility condition.
Theorem 1 leads to novel comparisons of Markov equilibria. In the complete information game, Markov equilibrium cannot sustain the Stackelberg payoff even when it is an equilibrium of the repeated game in general (Propositions 1 and 2). However, when the regime faces reputational incentives, the unique Markov equilibrium is the one where the long-run player plays their Stackelberg action at every period and perfectly pools with the committed type. [^3] This contrasts with past work, which shows that even under perfect monitoring there can be a continuum of equilibria, see Pei and Li (2021) Thus, the Markov refinement not only helps select the regime’s best equilibrium payoff, but allows the regime to strictly outperform their best Markov payoff in the complete information game when they have incentives for reputation formation.
I illustrate my results by returning to the chain store game. I show that when accommodating the entrant leads to some probability the game permanently ends, one of two things can happen depending on whether the credible beliefs restriction is adopted. First, in Markov equilibrium, there exists a unique equilibrium profile which secures the long-run player their Stackelberg payoff. However, in sequential equilibrium, the long-run player’s payoff may secure only their minimax payoff, exactly because of pathological equilibria where short-run players vary play in ways that are payoff-irrelevant. That the long-run player cannot secure a high payoff in sequential equilibrium contrasts sharply with the results of the KWMR chain store game absent exit. I interpret these results as giving qualified commitment foundations for the chain store game with exit.
Finally, I apply my model to repeated globalized regime change games with exit. A long-run player endogenously determines, at a cost, some coordination friction, after which short-run players see noisy signals about the regime’s strength and decide to attack. The game continues if and only if sufficiently few players attack that the regime is able to overcome the revolution in that period. Such a setup is similar to the game in Angeletos et al. (2007), but with reputational effects and endogenous regime actions. I use Theorem 1 to establish an asymptotic selection result, [^4] I consider the limit where both the idiosyncratic and aggregate noise vanish in order to satisfy my incentive compatibility condition. Theorem 2: In the unique Markov equilibrium, the regime attains their Stackelberg payoff if and only if the average strength of the regime is above the endogenously chosen risk-dominant level. Conversely, if the average strength is below the risk-dominant level, then the regime dies almost surely in the first period and attains an arbitrarily low payoff. Since observed equilibrium play jumps at the risk-dominant threshold, my result sheds light on why seemingly ex-ante similar regimes can differ dramatically ex-post in both (1) observed punishment and (2) length of survival.
The remainder of this paper is organized as follows. Section 1.1 reviews the related literature. Section 2 introduces the model. Section 3 establishes a folk theorem and a reputation payoff bound in Markov strategies, the main results of the paper. Section 4 applies the model to the chain-store game. Section 5 extends the framework to study repeated global games of regime change. Section 6 concludes. Omitted proofs are collected in the appendix. A supplemental appendix discusses the discounted repeated game.
1.1 Related Literature
This paper contributes to the extensive literature on when reputational incentives can select payoffs, started by Kreps and Wilson (1982) and Milgrom and Roberts (1982). I adopt the canonical framework of Fudenberg and Levine (1989) and Fudenberg and Levine (1992), who establish the generality of reputational effects in a wide class of infinitely repeated games with long-run and short-run players (see Fudenberg et al. (1990) and Fudenberg and Levine (1994)). There is a large and rich literature studying variations to their model; see Part IV of Mailath and Samuelson (2006) or Mertens et al. (2015) for a survey.
This paper is particularly related to Pei (2024) and Pei (2025), who study endogenous records as a key friction to reputation building. Importantly, Pei shows that endogenous records accord differing roles to the exit probability and patience of players, and consequently there is rich asymptotic behavior depending on which force in the discount rate is taken to $1$. In contrast, I endogenize the exit probability instead of records, so in particular allow it to always be bounded away from its patient limit. This leads to a different exercise than Pei’s models, and thus different predictions despite a similar decomposition of the exit probability. This paper also relates to Pei and Li (2021), who characterize equilibrium play in the Fudenberg and Levine (1989) model without exit; I characterize unique Markovian play with exit.
Relatedly, Abreu and Gul (2000) and the literature on reputational bargaining (see Fanning and Wolitzky (2022) for a review) assume the game ends when one player concedes, a particular form of exit. While the way endogenous exit is encoded is related, the models are quite different: for example, bargaining models are zero-sum, assume all players are long-run, and have very special exit structures. That these models are economically different can be seen in the new technical challenges that arise in my model and the distinctiveness of the qualitative conclusions that attain.
My paper also relates to models of “bad reputation” in extensive form reputation games (see Ely and Valimaki (2003), Ely et al. (2008), and Luo and Wolitzky (2025)) where a shutdown of statistical learning is similar to exit. I contribute to this literature by first directly working with the exit probability as a primitive instead of a derived object due to monitoring, and show that under this interpretation there are cases where reputation can once again secure a high payoff for the long-run player.
Using Markov refinements to make predictions in reputational repeated games is not new, and I see my refinement as contributing to a long and historical strand of literature that makes this assumption in order to say something meaningful about equilibrium play. For example, Pei (2016) and Lukyanov et al. (2025) use a Markov refinement to establish “inverse U-shaped” behavior in labor and goods markets, respectively, to reputational incentives, while Kostadinov and Roldán (2025) focus on Markov equilibrium to characterize the dynamics of optimal monetary policy with reputational concerns. Giannitsarou and Toxvaerd (2003) use Markov refinements to render the repeated global game (which is a leading application of my framework) tractable and make predictions. Bhaskar et al. (2013) provide a foundation for Markov perfection in long-run short-run games with bounded memory: they are the only purifiable equilibrium. In games with reputation, Sannikov and Faingold (2011) show that continuous-time games select a unique sequential equilibrium which is Markovian in the belief of the long-run player’s rationality—exactly my criteria. My paper gives an alternative justification for Markov criteria, as a minimal way of selecting payoffs in the repeated game with endogenous exit.
Finally, the regime change application builds on the rich literature on equilibrium selection in global games, started by Carlsson and van Damme (1993) and Morris and Shin (1998) (see Morris and Shin (2003) for a survey). The model’s key innovation in this area is to (1) allow the regime to choose, at a cost, the optimal punishment level, and (2) embed the game into a repeated regime change game reminiscent of Angeletos et al. (2007) and Chassang (2010) but with an i.i.d. state. The first change relates to a large literature in global games where some “large” player can take an action that modifies the global game and leads to new applied insights about regime change (see Angeletos et al. (2006), Edmond (2013), and Morris and Shadmehr (2024)). Proposition B.1 makes an independent contribution in this direction by establishing a novel comparative static that rationalizes why stronger or more patient regimes are more likely to be cruel, even when doing so harms the regime. The second change allows me to analyze reputational effects, and unlike Angeletos et al. (2007) and Angeletos and Pavan (2013), establish the existence of a unique Markov equilibrium.
2 Model
2.1 The Stage Game
There are $N+1$ players, $i\in[N]\cup\{L\}$, representing a long-run player $L$ who interacts in each period with $N$ short-run players. Each short-run player $i$ chooses an action from a compact set $a_{i}\in A_{i}$; let $a=\{a_{i}\}\in\prod_{i}A_{i}=A$ denote a profile of short-run player strategies. The long-run player chooses an action $c\in C\subset\mathbb{R}$, where $C$ is a countable (potentially finite) set including 0.555In the global games application, the set of actions will be a compact interval. Note this is essentially without loss of generality given Assumption 3: If $f:Q\to\mathbb{R}$ is continuous, where $Q\subset X$ is countable and dense, there is a unique continuous function $\overline{f}:X\to\mathbb{R}$ such that $\overline{f}|_{Q}=f$. However, the countability assumption helps significantly when defining equilibrium beliefs off-path, so I maintain it throughout the general analysis. Short-run player actions generate finite signal $Y$, drawn independently according to a distribution $\rho(\cdot|a,c)$. [^6] Short-run player monitoring is irrelevant to the analysis and included only for generality. The long-run player’s action is perfectly monitored.
The action profile $(a,c)$ jointly determines a survival probability, represented by a function $f:A\times C\to[0,1]$. This can be interpreted as the survival probability of a regime given a short-run action profile $a$ and long-run action $c$. The complementarity probability 1f(a,c)1-f(a,c) is the exit probability.
The long-run player’s expected stage payoff is thus given by $f(a,c)v(a,c)$, where $v(a,c)$ is the payoff they attain conditional on surviving. Short-run player $i$ has payoffs $u_{i}(a,c)$. Suppose all payoffs are bounded and continuous in all arguments. Moreover, assume that $f(a,c)<1-\xi$ for some $\xi>0$ always. This can be interpreted as saying there is always some (potentially arbitrarily small) probability exit occurs regardless of what actions are taken. [^7] This ensures the undiscounted payoff streams are well-defined, and in particular converge.
For any long-run player action $c$, denote by $\mathcal{N}(c)$ the Nash correspondence at $c$; that is,
$\mathcal{N}(c)=\left\{a:a_{i}\in\operatorname*{\arg\!\max}_{\overline{a}_{i}\in A_{i}}u_{i}(\overline{a}_{i},a_{-i},c)\right\}.$
Moreover, define
$a^{*}(c)=\operatorname*{\arg\!\min}_{a\in\mathcal{N}(c)}\left\{f(a,c)v(a,c)\right\}$
as the set of actions that minimize the long-run player’s stage-game payoff, subject to the fact short-run players are best-replying to each other given $c$ is the long-run player’s action. When there is no ambiguity, I also abuse notation and use $a^{*}(c)$ to mean a selection from the correspondence (for example, $f(a^{*}(c),c)$).
Throughout the analysis, it will be useful (expositionally) to impose the assumption a (myopic) long-run player who need not worry about exit has a dominant action, labeled at $c=0$, and that they benefit from commitment.
Assumption 1.
For all $a$, $0=\operatorname*{\arg\!\max}_{c\in C}f(a,c)v(a,c)=\operatorname*{\arg\!\max}_{c\in C}v(a,c)$. Moreover,
$0\not\in\operatorname*{\arg\!\max}_{c\in C}\min_{a\in A}\left\{f(a,c)v(a,c)\right\}$
The first part of the assumption implies the set of payoffs attainable in the one-shot game can be represented as $\{f(a,0)v(a,0)\}_{a\in\mathcal{N}(0)}$, i.e. the profiles where the long-run player chooses 0 and short-run players best reply. This simplifies the exposition by simplifying the “worst” punishments from deviating from a prescribed equilibrium (though at the cost of substantially more notational overhead, is not necessary). Towards this end, let $\overline{V}_{0}=\max_{a\in\mathcal{N}(0)}\frac{f(a,0)v(a,0)}{1-f(a,0)}$ denote the long-run player’s largest payoff from playing the static optimum repeatedly, and analogously let $\underline{V}_{0}$ be their smallest such payoff. The second part of the assumption guarantees guarantees the long-run player strictly benefits from the ability to build a reputation when compared to their worst stage-game Nash payoff.
2.2 The Repeated Game
The long-run player plays the stage game in each of infinitely many periods ($t=0,1,2,\dots$), weighting each period equally. In each period, they face a population of short-run players who take myopic best replies.
Time-$t$ histories are elements $h^{t}=(y_{s},c_{s})_{s=0}^{t-1}$ of past short-run player signals and long-run player actions. Let $H^{t}$ denote the set of time-$t$ histories, $H=\bigcup_{t}H_{t}$ the set of all histories, and $H^{\infty}=(Y\times C)^{\infty}$ the set of all infinite histories. Short-run player strategies are functions $\sigma_{i}:H\to\Delta(A_{i})$, while the (rational) long-run player’s strategy is a function $\sigma_{L}:H\to\Delta(C)$. [^8] Following Mertens et al. (2015), there is a canonical mapping from strategies $\sigma:H\to\Delta(\mathcal{X})$ and strategies $\sigma:H^{\infty}\to\Delta(\mathcal{X})$ for $\mathcal{X}\in\{C,A\}$. When it is useful (i.e. when topologizing $\sigma$), I will use these two codomains interchangeably.
Given a sequence of realized actions $(a_{t},c_{t})_{t=0}^{\infty}$, the long-run player’s payoff is
$\mathcal{V}(\{a_{t},c_{t}\})=\sum_{t=0}^{\infty}\left(v(a_{t},c_{t})\prod_{s=0}^{t}f(a_{s},c_{s})\right).$
Note payoffs here are not discounted. SUPPLEMENTAL APPENDIX studies the discounted repeated game and shows the game here is the standard limit as players become patient, but with exit. However, discounting payoffs does not add anything to the analysis and, given the exit structure, would in fact substantially complicate the analysis. Consequently, I omit the discount rate $\delta$ in the main body of the paper, and refer to this as the case of “extreme patience” for the long-run player.
Define the long-run player’s lower Stackelberg strategy as
$V_{*}=\sup_{c\in C}\min_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}v(a,c)=\sup_{c\in C}\min_{a\in\mathcal{N}(c)}\frac{f(a,c)v(a,c)}{1-f(a,c)}$
and let $c^{*}$ be a maximizer in $V_{*}$ (throughout, suppose such a maximizer exists). Let $V^{*}$ be the upper Stackelberg payoff, defined analogously.
Payoffs are written such that exit is permanent: once exit has occurred, the long-run player leaves the game, and hence their payoff is zero in all future periods regardless of any hypothetical actions that would have been taken.
The long-run player can either be a rational type, $\omega_{R}\in\Omega$, who seeks to maximize their discounted payoff, or one of potentially many committed types, $\omega_{c}\in\Omega$, who plays $c$ at every possible history. Short-run players hold a full-support prior belief $\mu_{0}\in\Delta(\Omega)$ over the long-run player’s type. Write $\sigma(\cdot,\omega_{c})\equiv c$ for the strategy of a commited type $\omega_{c}$.
For any strategy $\sigma$, define an associated belief system $\mu:H\to\Delta(\Omega)$ which outputs some belief about the long-run players’ type for all histories. Together, $(\sigma,\mu)$ is an assessment. Given an assessment, let $\overline{\sigma}_{L}(h^{t})$ denote the unconditional strategy of the long-run player averaging over $\mu(h^{t})$, and $\sigma_{L}(h^{t};\omega)$ denote the strategy of a type $\omega$ long-run player. Analogously, let $\overline{\mathbb{P}}^{\sigma}$ denote the probability measure over histories generated by the unconditional strategy $\sigma$, and $\mathbb{P}^{\sigma}$ the measure generated when $\omega=\omega_{R}$. The main results of this paper study the set of (rational) long-run player payoffs that can be attained in sequential equilibria $(\sigma,\mu)$, defined as follows:
Definition 1.
An assessment $(\sigma,\mu)$ is a sequential equilibrium if:
  1. (1)
    Sequential Rationality. At every history $h^{t}$, the continuation strategy $\sigma_{L}(\cdot|h^{t})$ maximizes the long-run player’s continuation payoff given $\{\sigma_{i}(h^{t})\}$ and $\mu(\cdot|h^{t})$.
  2. (2)
    Short-Run Best Replies. At every history $h^{t}$, $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(\overline{\sigma}_{L}(h^{t}))$.
  3. (3)
    Bayesian Updating. At every history $h^{t}\in\text{supp}(\overline{\mathbb{P}}^{\sigma})$, $\mu(h^{t})=\mu(\cdot|h^{t},\sigma)$, i.e. the conditional belief over $\Omega$ given history $h^{t}$ and strategy $\sigma$.
  4. (4)
    Consistency. There exists a sequence $(\sigma^{n},\mu^{n})$, with $(\sigma^{n},\mu^{n})\to(\sigma,\mu)$ such that
    1. (a)
      $\text{supp}(\overline{\mathbb{P}}^{\sigma^{n}})=H$ for all $n$, and $\mu(\cdot|h^{t},\sigma^{n})=\mu^{n}(h^{t})\in\Delta(\Omega)^{o}$, and
    2. (b)
      $\sigma_{L}^{n}(\cdot,\omega_{c})=c$ at every history $h^{t}$.
An assessment $(\sigma,\mu)$ is a Markov perfect equilibrium if it is a sequential equilibrium and moreover at any histories $h^{t},h^{s}$,
$\mu(h^{t})=\mu(h^{s})\implies\sigma(h^{t})=\sigma(h^{s})$
Because $\sigma:H\to\Delta(C)$ maps an infinite set into another, the notion of convergence is not immediately clear. Say $(\sigma^{n},\mu^{n})$ converges to $(\sigma,\mu)$ if $\sigma^{n}\to\sigma$ converges in the strong topology (see Appendix B of Luo (2025)) and $\mu^{n}\to\mu$ converges weakly. At any history $h^{t}$ where $\mu(h^{t})$ is degenerate, sequential equilibria are equivalent to subgame perfect equilibria. Consistency of equilibrium profiles will be useful when evaluating the value of deviations the long-run player may consider which are off the path of play (and not supported by some commitment type); absent some consistency-type condition, the value of such deviations would be hard to define in a principled way. Note the “trembles” defined here for consistency do not change the prior belief over $\Omega$ (which would then change $\mu^{n}$). A standard compactness argument shows this is without loss of generality (one could add the requirement $\mu_{0}^{n}\to\mu_{0}$ as well), but for simplicity I omit this additional notation.
Markov perfection requires the stronger assumption that at any two histories where short-run player posterior beliefs about long-run play are the same, players take the same actions. [^9] It may seem unnatural to require the long-run player take the same action when short-run player beliefs are the same. The definition of Markov perfection can be relaxed to require only that short-run players strategies are measurable in $\mu$; however, since there is a single long-run player, their strategy can then be taken to be (without loss of generality) also measurable in $\mu$, less a public randomization device. While such a refinement may seem a-priori restrictive, I show it arises naturally when studying the effect exit has on reputational incentives (see Theorem 1).
Let $\underline{U}$ and $\overline{U}$ be the long-run player’s lowest and highest payoff from any sequential equilibrium in the game. [^10] There is some potential ambiguity about whether sequential equilibrium need to exist when $C$ is infinite (see Myerson and Reny (2020) for some discussion and general fixes). However, given $\Omega$ is finite and $C$ is countable, the problem is substantially simpler—I show in Proposition 1 and Theorem 1 that sequential equilibria as defined above exist, so in particular $\underline{U},\overline{U}$ are finite. Similarly, let $\underline{U}^{M}$ and $\overline{U}^{M}$ be the long-run player’s lowest and highest payoff from any MPE in the game.
3 Commitment Payoff Theorems
This section contains the main results of this paper, which characterize the payoff set in the complete information game with and without reputation, in sequential and Markovian equilibria. Informally, I establish the following table, which summarizes the takeaway from the results in this section. Together, these results form the main contribution of this paper.
non-Markov Markov
Reputation Anti-Commitment Theorem (Prop. 3) Stackelberg Payoff (Theorem 1)
No Reputation Folk Theorem (Prop. 1) Static Nash Payoff (Prop. 2)
3.1 Complete Information
Consider the complete information game where there is no uncertainty about the long-run player’s rationality. Say a folk theorem holds in the game if the set of payoffs attainable as part of some equilibrium in the complete-information game is the set of feasible and individually rational payoffs for the long-run player (here the set $V_{0},V^{*}]$), supposing short-run players’ myopically best reply. Proposition 1 characterizes exactly when a folk theorem holds in our game. Arguments from Fudenberg and Levine (1994), Fudenberg et al. (1994) (which would immediately imply a folk theorem) do not necessarily apply because exit can “loom large,” and thus lead to periods where the effective discount rate is small even without discounting, i.e. $\delta=1$. Thus, there exist games where not all feasible, individually rational payoffs are attainable. Proposition 1 characterizes, as a function of long-run player payoffs, exactly when a folk-like theorem holds as a function of the long-run player’s payoff, and formalizes the sense in which efficient payoffs can be sustained only when the future “looms large” given the exit structure. Let $L(a,c)=\frac{f(a,c)}{f(a,0)}$ be the likelihood ratio of exit when the long-run player chooses $c$ at some action instead of 0 given short-run player action $a$.
Proposition 1.
Let $\mu_{0}(\omega_{R})=1$ and let $(a^{*},c^{*})$ attain $V^{*}$. Then
$\underline{U},\overline{U}]=[\underline{V}_{0},V^{*}]$
if and only if for any $(a^{*},c^{*})$ attaining $V^{*}$,
$(V^{*})L(a^{*},c^{*})-\underline{V}_{0}\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*}).$
If either of the two (equivalent) statements above holds, say the game admits a folk theorem. Proposition 1 gives a good baseline for when all feasible payoffs can be obtained in some equilibrium; moreover, the relevant incentive compatibility in Proposition 1 will give, under some monotonicity and uniqueness conditions, an exact characterization of when the long-run player exactly secures their Stackelberg payoff (Corollary 1). The converse direction of the result implies the long-run player’s best equilibrium payoff is bounded away from their Stackelberg payoff even as players get extremely patient (i.e. the folk theorem fails) so long as the incentive compatibility condition fails.
The proof is relatively straightforward and can be found in the PROOF OF PROPOSITION 1 The key step in the proof is to write the incentive constraint for the long-run player, supposing they obtain $V^{*}$ if they play $c^{*}$ and $\underline{V}_{0}$ otherwise and algebraically rearrange terms to conclude it is equivalent to the condition in the proposition. This then implies playing $c^{*}$ is self-enforcing if (and only if) the condition in the result is satisfied.
Importantly, though, the equilibria constructed in the proof of Proposition 1 to sustain a high payoff for the long-run player require that punishments off the equilibrium path are non-Markov, in the sense that they are not measurable in $\mu_{t}(h^{t})$ (which is constant in the complete information case). Given this observation, it is perhaps unsurprisingly that the Markov restriction can strongly refine the set of payoffs: in the complete information game, it does so exactly by holding the long-run player down to their exact static payoff.
Proposition 2.
Let $\mu_{0}(\omega_{R})=1$. Then $\underline{U}^{M},\overline{U}^{M}]=[\underline{V}_{0},\overline{V}_{0}]$.
This property of the complete-information Markov game will be important in disciplining the value of deviations in the game with reputational concerns. The proof follows immediately once it becomes clear that the Markov refinement requires the continuation strategies to be independent of the history (since $\mu(h^{t})=1$ at all histories in all sequential equilibria). For completeness, I give a formal argument in the PROOF OF PROPOSITION 2 . Note that Proposition 2 applies to any repeated game, not necessarily just those with exit.
3.2 Guaranteeing Stackelberg Payoffs
What classes of reputation results might one hope for in the game with exit? Recall that classical reputation results hold only after many111111“Many” here depends formally on the speed of learning and the prior belief of rationality; see Gossner (2011) for a more formal discussion. (finite) periods have passed, as the long-run player must teach their opponents that they are playing the commitment action. Yet in the game with exit, they may not have the luxury of time; the game may have ended far before short-run player beliefs about long-run play have concentrated on $c^{*}$, keeping the long-run player from securing a high payoff.
Thus exit perturbs the long-run player’s strategic incentives in two ways: first, by making the future less valuable (and hence the promise of future payoffs less motivating to discipline signaling behavior today), and second, by shortening the length of time they have to convince short-run player beliefs about their (committed) play.
The first problem, as evidenced by Proposition 1, can be resolved by adding an appropriate incentive compatibility constraint for the long-run player. The second problem, however, is not so easy to fix: indeed, I will need the full force of the Markov refinement in order to iron out beliefs which are adversarial to guaranteeing the long-run player their Stackelberg payoff. The technical contribution of Theorem 1 is that these two ingredients exactly characterize when one can obtain a commitment payoff theorem in the game with exit.
Definition 2.
Say the long-run player satisfies (weak) $c^{*}$ incentive compatibility if for each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$(V_{*})L(a,c^{*})-\overline{V}_{0}>(\geq)\left(v(a,0)-v(a,c^{*})L(a,c^{*})\right)$
This incentive compatibility condition is stronger than the one found in the folk theorem, Proposition 1. Intuitively, it requires that for all potential short-run player best replies $a$ that can arise on the equilibrium path, the future looms sufficiently large that playing $c^{*}$ is enforceable against playing 0 when the long-run player can secure their Stackelberg payoff from $c^{*}$ and static Nash otherwise. In the leading application, where the likelihood ratio and stage-game payoffs depend only on $c$ (i.e. $L(a,c^{*})=L(c^{*})$ and $v(a,c)=v(c)$), this condition boils down exactly to the incentive compatibility condition in Proposition 1 (see Corollary 1). For now, it delivers the following theorem, (modulo one last definition for the converse).
Definition 3.
Primitives $(f,v,u)$ are monotone if $A$ can be ordered such that for each fixed $c$, $f(a,c)$ is strictly decreasing in $A$, $v(a,c)$ is nonincreasing, and $\mathcal{N}(\alpha c)$ is decreasing (in the strong set order) in $\alpha$.
Theorem 1 (Markov Commitment).
Let $\Omega=\{0,c^{*}\}$ with $\mu_{0}(\omega_{R})<1$. If the long-run player satisfies $c^{*}$ incentive compatibility, then in all Markov equilibrium $\sigma_{L}(h^{t})\equiv c^{*}$. In particular, $\underline{U}^{M}=V_{*}$.
Conversely, if $(f,v,u)$ is monotone and the long-run player does not satisfy weak $c^{*}$ incentive compatibility, then
$\limsup_{\mu_{0}(\omega_{R})\to 1}|\underline{U}^{M}-V_{*}|>0.$
Proof.
The forward direction. Suppose the long-run player satisfies $c^{*}$ incentive compatibility. We show in all Markov perfect equilibria the long-run player plays their commitment action $c^{*}$ and short-run players best respond, securing the long-run player $V_{*}$. To do so, we establish a number of lemmas whose (technical) proofs are relegated to the PROOF OF LEMMA 1
First, we must discipline the inferences and payoffs that short-run players have when the long-run player does not imitate the commitment type.
Lemma 1.
Let $(\sigma,\mu)$ be a sequential equilibrium. For all $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$, $\mu(h^{t})(\omega_{R})=1$.
Lemma 1 (along with Proposition 2) guarantees that, off the equilibrium path, the highest payoff the long-run player can obtain is $\overline{V}_{0}$. Our next lemma disciplines the set of actions the long-run player can take in any equilibrium.
Lemma 2.
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $\text{supp}(\sigma_{L}(h^{t}))\in\{0,c^{*}\}$.
Lemma 2, along with Lemma 1 are enough to imply (following an argument similar to the first part of the proof of Proposition 1 and taking $\alpha=1$), that always playing $c^{*}$ is a Markov perfect equilibrium.
Our last step is to show that this is the only Markov perfect equilibrium. Clearly, always playing 0 cannot be Markov perfect under assumption. To see why, take $\alpha=0$ and fix any $a\in\mathcal{N}(0)$. Rearranging the incentive compatibility constraint implies
$f(a,0)[v(a,0)+\overline{V}_{0}]<f(a,c^{*})[v(a,c^{*})+V_{*}]$
so the long-run player can strictly profitably deviate by always choosing $c^{*}$ in any period they were supposed to deterministically play 0. Finally, we show that the long-run player prefers not to mix in Markov equilibria.
Lemma 3.
Let $(\sigma,\mu)$ be Markov perfect. Then for all $h^{t}$, $|\text{supp}(\sigma_{L}(h^{t}))|=1$.
The intuition is that in order to sustain mixing, the long-run player must be indifferent between their upper Nash and lower Stackelberg payoffs, modulo the cost today. When this difference is large and exit does not shut the game down early (the second condition), then this can never be case, which ensures the long-run player is never indifferent between actions.
The proof of the converse is constructive and can be found in the PROOF OF THEOREM 1 (Converse) The basic idea is that, at any $\alpha$ where the incentive compatibility constraint holds with equality, we find an equilibrium where the long-run player mixes at the time $t=0$ public history. Monotonicity of the game then ensures that at mixing histories, they underperform their commitment payoff. The condition that $\mu_{0}$ is small rules out equilibria where $\mu_{0}(\omega_{c^{*}})\approx 1$ and the long-run player can secure a very high payoff in the first period even if they play 0. ∎
The Markov refinement here disciplines short-run player learning; long-run players, in a Markov equilibrium, begin teaching short-run players about play at time $t=0$ (as opposed to needing to slowly form their reputation). Reputation thus only has a “backwards looking” component: instead of needing to teach short-run players about future play by taking some commitment action today, they need only maintain their current reputation that they will continue to do what they have already done. Theorem 1 therefore can be interpreted as saying in reputational games with exit that this backwards-looking force is the economically relevant one, i.e. the one that is not blocked by the exit structure.
Monotonicity of the game is satisfied in natural applications, including the coordination game later studied in this paper. For example, it is satisfied whenever the exit probability increases with “more aggressive’ actions $a\in A$, $v$ depends only on the long-run player’s action, and short-run players play a strictly submodular game with unique best replies (as is the case in the globalized coordination game analyzed later) in $(a,c)$ so that higher levels of long-run player actions lead to less aggressive sets of actions. Monotonicity is sufficient to imply that when the long-run player is indifferent between several actions, they cannot secure a higher payoff than their pure Stackelberg payoff. This rules out situations where the long-run player benefits from mixed Nash equilibrium. In particular, it ensures that when mixing occurs in the first period (which is used to prove the converse to Theorem 1), the long-run player must strictly underperform their commitment payoff in such an equilibrium.
In general, the incentive compatibility condition given in the theorem can be somewhat unwieldy to interpret. However, it turns out that there is a special subclass of problems in which the ability to attain the Stackelberg payoff (uniquely) is equivalent to asking if it possible to support that payoff in the complete-information game at all.
Corollary 1.
Suppose $v(a,c)=v(c)$, $L(a,c)=\mathcal{L}(c)$ for some $\mathcal{L}(c)$, $|\mathcal{N}(0)|=1$, and $V^{*}=V_{*}$. Then the following are equivalent.
  1. (1)
    There exists $\overline{\mu}_{0}>0$ such that for all $\mu_{0}<\overline{\mu}_{0}$, $\delta>\overline{\delta}$, $\underline{U}^{M}=V^{*}$.
  2. (2)
    The game admits a folk theorem.
The proof simply verifies that, under the stated conditions, the incentive compatibility conditions of Proposition 1 and Theorem 1 coincide. It is delegated to the Proof of Corollary 1
[…]
3.3 Non-Markov equilibrium
Theorem 1 focuses on Markov strategies and leaves open the question of when reputation can help refine payoffs away from the Markov case. Without putting additional restrictions on the survival probability function $f$, is it possible to say anything about when the long-run player can establish a reputation outside of Markov strategies? I next show the answer is no—even with $c^{*}$ incentive compatibility, one cannot refine the payoff set in general, as the long-run player’s payoff may be arbitrarily low.
Proposition 3 (Anti-Commitment.).
Fix payoffs $(u,v)$ such that there exists $c$ where
$\min_{a\in\mathcal{N}(c)}v(a,c)>0$. Then for any $\varepsilon>0,M>0$, there exists $f$ such that
  1. (1)
    The game admits a folk theorem and $V^{*}\geq M$.
  2. (2)
    $\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq\varepsilon$.
The first assumption is a nontriviality condition guaranteeing the existence of a non-Nash commitment posture, conditional on a high probability of survival. The limit in the second statement guarantees the probability of commitment is not “so high” that what the rational long-run player does will not affect their payoffs (i.e. if $\mu_{0}(\omega_{c^{*}})\approx 1$).
Proposition 3 is stronger than the simple statement that the long-run player’s payoff can be arbitrarily low, which can be achieved by simply setting $f(a,c)\leq\eta$ for all pairs $(a,c)$ and arbitrarily low $\eta$. In particular, the first statement guarantees that the long-run player attains a low payoff even if (1) the folk theorem holds, so they can attain their Stackelberg payoff, and (2) such a Stackelberg payoff is arbitrarily high. In this way, Proposition 3 guarantees that in sequential equilibrium, even in the prescence of reputational incentives, the payoff set can become large absent strong assumptions (stronger than $c^{*}$ incentive compatibility) on the exit function. One consequence of Proposition 3 is that as $\mu_{0}(\omega_{R})\to 1$, the set of sequential equilibria payoffs with reputation converges to the set of sequential equilibria payoffs without, in stark contrast to the baseline payoff selection results of Fudenberg and Levine (1989) without exit.
Proposition 3 highlights the two important roles of Markov equilibrium in establishing the positive reputation result of Theorem 1. First, it disciplines the feasible payoff set after rationality is revealed (thus decreasing incentives for “revealing rationality”), making it hard for short-run players to reward tomorrow actions which are not the commitment action. Second, it ensures short-run players believe the commitment action will be played whenever reputation foundations have bite, which ensures that exit does not “loom large” in the first period (c.f. the discussion succeeding Theorem 1. Markov beliefs thus rule out (potentially pathological) equilibria where the long-run player is rewarded in the future in a complicated way for revealing they are rational today, even though doing so will cause the game to end with arbitrarily high probability today. The proof is relegated to the PROOF OF PROPOSITION 3
4 The Chain Store Game
Consider the canonical chain store game, with two variations: first, suppose the game is infinitely repeated, so the monopolist must continually face entrants over and over again, and second, suppose player actions can lead to exit in the game. The first assumptions implies that, without exit, the results of Fudenberg and Levine (1989) apply, and hence the long-run player can secure their Stackelberg payoff in all Nash equilibria as they become extremely patient. The goal of this section is to understand how exit perturbs these results. In incumbent-entrant games, exit may be a particularly appropriate assumption: accommodating an entrant can lead to the incumbent being displaced form their position, while fighting may increase the probability they remain an incumbent even after entry. This can lead to an additional channel for and challenges to reputation formation.
Formally, suppose in each period the long-run player ($I$) chooses to fight or accomodate an entrant, but only after the entrant ($E$) has decided whether or not to enter. Let stage game payoffs be as follows, where the first entry is the long-run player’s payoff. [^12] These specific payoffs for the chain store game are due to Harry Pei’s Lecture Notes. Interpret here “accomodate” as the action 0.
[…]
Figure 1: Payoffs in the Chain Store Game
Moreover, define $f(a,c)$ to take the values in 1, where $\varphi,\varepsilon,\eta$ are strictly positive and $1-\eta\geq\max\{\varphi,\varepsilon\}$, so survival occurs with higher probability if the entrant chooses Out.
Fight Accomodate
In $\varphi$ $\varepsilon$
Out $1-\eta$ $1-\eta$
Table 1: Exit Probabilities for the Chain Store Game
Note that these payoffs imply the long-run player’s commitment action $c^{*}$ is $F$ for all $\delta$, but that accomodate is the unique Nash equilibrium. These observations (in addition to ranking $F\succ A=0$) imply that Assumption 1 is satisfied. The chain store game admits the following result.
Proposition 4.
Fix $\varphi>0$. There exists $\overline{\varepsilon}>0$ such that for all (and only) $\varepsilon<\overline{\varepsilon}$
$\limsup_{\mu_{0}(\omega_{R})\to 1}|\underline{U}^{M}-V_{*}|=0.$
Moreover, there exists $e(\cdot)$, vanishing at 0 such that
$\limsup_{\mu_{0}(\omega_{R})\to 1}\underline{U}_{0}\leq e(\varepsilon).$
The formal proof is in the PROOF OF PROPOSITION 4 The intuition behind the first part is to show that as $\varepsilon$ vanishes, payoffs satisfy $c^{*}$-incentive compatibility; since the game is monotone in the order $In\succ Out$, this implies the second statement. The second part follows by taking $\varepsilon$ small and constructing a revelation equilibrium as in the proof of Proposition 3.
Three observations are in order. First is that in Proposition 4, the key parameter was the probability of exit given that the incumbent entered the game. Intuitively, varying $\varepsilon$ varies the effective discount rate but also the incentives to take different actions; lower $\varepsilon$ increase the “cost” of Nash equilibrium and make reputation formation more attractive to the long-run player, especially relative to fighting. Mathematically, this is evidenced through the likelihood ratio $\frac{\varphi}{\varepsilon}$; as $\varepsilon\to 0$, the likelihood ratio of survival after fighting explodes relative to that after accommodating. Second is the contrast of Proposition 4 with known results from reputation a la Fudenberg and Levine (1989). In particular, the second part of the proposition highlights how reputation can fail with exit even with extremely patient players and high likelihood ratios for survival conditional on the commitment action. This shows that the Markovian refinement is critical for any hope of selecting a unique equilibrium, exactly because it rules out nonstationary equilibrium where the exit cost becomes “sunk” to the long-run player, who chooses their stage-game action conditional on the fact they have already survived today. Finally, I am able to give a unique behavioral prediction in Markov strategies in this game in which the incumbent always fights and short-run players never enter (i.e. deterrence is successful), unlike the indeterminancy of behavioral predictions in reputation games in general (c.f. Pei and Li (2021)).
5 Regime Change
5.1 The Global Games Model
Consider the problem faced by an authoritarian government (player $L$) who must decide how they should punish dissidents who engage in a revolution. In each period, having successfully defended against an attack, the government can choose to punish dissenters by decreasing their per-period utility by c-c at a cost of $\kappa(c)$, where $\kappa$ is a continuous function satisfying $\kappa(0)=0$. Let $C=[0,\overline{c}]$ and moreover suppose $\kappa(\overline{c})>1$, so that the government never chooses the maximally feasible value of punishment. [^13] Note $C=[0,\overline{c}]$ is distinct from the assumption made in the main model that $C$ is countable. This is useful in defining $c^{*}(0)$, after which it is sufficient to assume $C=\{0,c^{*}(0)\}$ for the remainder of the application. [^14] This is an Inada-type condition that ensures the government’s problem is always well-defined
The (unit mass, continuum) dissidents now face the following problem. In each period, they know the government is able to fend off $\theta$ attackers, where $\theta\sim P_{\theta}$ is drawn independently in each period, with $\mathbb{E}_{P_{\theta}}[\theta]=\mu$ and $\text{var}(\theta)=\tau^{2}$.  [^15] In practice, I will consider the case where $\tau^{2}$ is particularly small. This can be interpreted as a world where the regime has strength $\mu$ and experiences a small idiosyncratic shock in each period to that strength. Each dissident $i$ cannot see $\theta$ perfectly, instead observing an imperfect signal $x_{i}$ of $\theta$ with some idiosyncratic noise, $x_{i}=\theta+\varsigma\varepsilon_{i}$, where $\varepsilon_{i}\sim F_{\varepsilon}$ is i.i.d. across each player. Suppose players’ information sets are sufficiently well-behaved that they satisfy the following regularity conditions:
Assumption 2 (Informational Regularity).
The following hold.
  1. (1)
    $F_{\varepsilon}$ and $P_{\theta}$ admit strictly positive, bounded continuous densities; $F_{\varepsilon}$ is symmetric around 0, and $P_{\theta}$ is symmetric around $\mu\in(0,1)$.
  2. (2)
    The conditional distribution $F_{\theta}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
  3. (3)
    The conditional distribution $F_{x}(\cdot|x)$ has strictly positive density and is decreasing in its second argument.
Here, $F_{\theta}(\cdot|x)$ is the first order belief over fundamentals of players supposing they see signal $x$ and have prior $P_{\theta}$; $F_{x}(\cdot|x)$ is the second-order conditional belief about other players’ signals (i.e. player $i$’s thought about others’ signals after receiving $x$). The symmetric assumption on $F_{\varepsilon}$ allows it to be interpreted as white noise. I require $\mu\in(0,1)$ so that the state is not in a dominance region.
Assumption 2 also guarantees $F_{x}$ admits a strictly positive density and that $F_{x}(x|\theta)$ is decreasing in the realized state; moreover, $F_{\theta}(\cdot|x)$ and $F_{x}(\cdot|x)$ are differentiable as the noise term is additive. These smoothness assumptions ensure a first order approach for players is valid, and generalize known conditions standard in the global games literature. [^16] I am not aware of a global games setting studied in the literature that does not satisfy these assumptions; for example it nests normal-normal information structures.
Having seen their private signals, the dissidents independently and simultaneously choose an action $a_{i}\in\{0,1\}$ in each period, with $a=A=\int a_{i}di$ equal to the total mass of agents taking the aggressive action (i.e. revolt), $a_{i}=1$. If enough dissidents defect (i.e. $A\geq\theta$), then the revolt is successful, the regime is overthrown, and the game ends. The continuation probability $f(A,c)$ is exactly the probability the revolt is successful, is given by $f(A,c)=\mathbb{P}_{\theta}(A\geq\theta)$, and is independent of $c$.
Players’ payoffs are given by
[…]
Note here payoffs mirror standard global games payoffs (see i.e. Frankel et al. (2003)) where $\theta$ is the peg of the regime, but with the cost of failed coordination (i.e. the coordination friction) flexibly set by the regime at some cost. Standard arguments from the global games literature imply the following.
Proposition 5.
Suppose $\varsigma<\tilde{\varsigma}$ for some $\tilde{\varsigma}>0$ sufficiently small. Then for all $c$, there exists functions $x^{*},\theta^{*}:\mathbb{R}_{+}\to\mathbb{R}$ such that
  1. (1)
    The strategy $s(x,c)=\mathbf{1}\left\{x<x^{*}(c)\right\}$ is the (essentially) unique strategy surviving iterative deletion of strictly dominated strategies.
  2. (2)
    $x^{*}(c)\in\mathcal{C}^{1}$, and is strictly decreasing in $c$.
  3. (3)
    $\theta^{*}(c)\in\mathcal{C}^{1}$, is strictly decreasing in $c$, and $\int s(x,c)dF_{x|\theta}\geq\theta$ if and only if $\theta\leq\theta^{*}(c)$.
Proposition 5 implies not only that $\mathcal{N}(c)$ is single-valued, but that both the set of states which successfully guard against attacks and the cutoff state are themselves monotone at a cutoff, whose evolution is well-behaved. These properties will be useful in the proof of Theorem 2 and is proven in the PROOF OF PROPOSITION 5 Note Assumption 1 is satisfied as well as $1-\kappa(c)<1-\kappa(0)$ for all $c>0$, and the exit probability depends only on $A$.
Finally, the game is clearly monotone when $A$ is ordered in the standard way in $\mathbb{R}$. Hence the globalized game satisfies the assumptions of the general model.
5.2 Selecting Equilibria
When is it that the threat of punishment is enough to deter revolution? Since the game is monotone and satisfies the conditions of Corollary 1, $c^{*}$-incentive compatibility is both necessary and sufficient for the regime to deter revolution (as much as possible) in the unique Markov equilibrium. Rewriting the incentive constraints to the simpler setting of the global game described above yields that $c^{*}$-incentive compatibility is equivalent to the inequality
$V_{*}-\underline{V}_{0}>\kappa(c^{*}).$
In the global games setting, $c^{*}$ incentive compatibility is related to the degree of dispersion of higher order beliefs around the risk-dominant threshold at the average, i.e. whether $\mu>(<)\frac{1}{1+c^{*}}$. Here, recall value $\frac{1}{1+c^{*}}$ is the risk-dominant signal threshold selected by the global game as signals become arbitrarily precise, i.e. $\lim\limits_{\varsigma\to 0}x^{*}(c^{*})=\frac{1}{1+c^{*}}$ by Frankel et al. (2003).
What is the right commitment posture $c^{*}$ to consider when comparing to $\mu$? There are two concerns. First, since $c^{*}(\varsigma)$ varies with the noise precision, the risk-dominant selection threshold varies with the noise precision as well. Second, the underlying probability distributions $(F_{\varepsilon},P_{\theta})$ will also affect the optimal choice for the regime and affect their commitment payoff. To resolve this problem, consider the set
$c^{*}(0)=\limsup_{\varsigma\to 0}c^{*}(\delta,\varsigma)$
where $\limsup A_{n}$ are the subsequential limits of all selections from $A_{n}$. That $c^{*}(0)$ is nonempty follows from a compactness arguments, noting $c^{*}(\varsigma)$ is well-defined and nonempty for each $\varsigma$ and hence so must $c^{*}(0)$. Throughout, I will fix some $c^{*}\in c^{*}(0)$ and appeal to continuity of the problem in parameters to make the argument that the same statement holds for $c^{*}(\varsigma)$.
Throughout the statement of Theorem 2, suppose $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ for some $c^{*}\in c^{*}(0)$. I also assume by fiat that $c^{*}\neq 0$, which is assumed throughout the main exposition, so that the commitment type is distinguishable from one playing static Nash. [^17] This assumption can be derived by putting mild regularity conditions on $\kappa$, for example that $\kappa^{\prime}(0)=0$.
Theorem 2.
For all $\varepsilon>0$, there exists $\overline{\tau}>0$ and $\overline{\varsigma}(\tau)$ vanishing as $\tau\to 0$, such that if $\text{var}(\theta)=\tau<\overline{\tau}$, and $\varsigma<\overline{\varsigma}(\tau)$, then
  1. (1)
    If $\frac{1}{1+c^{*}}<\mu$, $\underline{U}_{0}^{M}=V^{*}$ and $f(a^{*}(c^{*}))>1-\varepsilon$.
  2. (2)
    If instead $\frac{1}{1+c^{*}}>\mu$, the $\max\{\underline{U}_{0},\overline{U}^{M}\}<\varepsilon$.
Theorem 2 highlights a stark difference between the regime’s Stackelberg payoff depending on where their average regime strength when (1) the average becomes an arbitrarily precise summary of the state in each period, (2) the regime becomes extremely patient, and finally (3) the global games perturbation vanishes. Importantly, it delivers a novel implication of risk dominance that is absent from the static game: the regime can only successfully deter rebellion if and only if they are, on average, able to deter rebellion at the risk-dominant level. Should they not be able to do this, then exit immediately happens with high probability.
The result stands in contrast to two seminal papers in the literature on dynamic global games with exit. First, it augments the model of Angeletos et al. (2007) to (1) small i.i.d. shocks to the state $\mu$ and (2) reputational incentives for the regime and shows how these perturbations can select a version of the “survival dominant” equilibrium (i.e. survival only if $\mu>\frac{1}{1+c^{*}}$) as the unique Markov equilibrium. Finally it augments the reputational global game model of Huang (2017) to allow for small i.i.d. shocks and a rich exit process, and shows how the regime can fail if their average strength is too weak. This stands in stark contrast to his main result, where there are no attacks that ever occur on the equilibrium path. The differences in these results suggest that taking exit incentives seriously in reputational models can lead to rich results which are more indicative of real-world phenomena, where successful attacks do happen if speculator’s sense the regime is “likely weak.”
The intuition behind the forward direction of Theorem 2 follows from Theorem 1 and a standard concentration inequality argument: if $\mu>\frac{1}{1+c^{*}}$, then the cutoff state as the noise vanishes is below the mean, so as the variance vanishes the regime must survive most of the time. Conversely, if $\mu<\frac{1}{1+c^{*}}$, then most of the time the regime dies, and hence they must attain a low payoff. A formal proof can be found in the PROOF OF THEOREM 2
6 Discussion
This paper studied a model of reputation where players’ actions endogenously affected the probability of exit in the game (and hence the effective discount factor was endogenous). Such endogeneity creates several difficulties for the analysis of the repeated game. First, folk theorems need no longer hold because the future can be “small” relative to actions today, making enforceability infeasible. Second, the addition of reputation effects need not select the long-run player’s Stackelberg payoff even when the long-run player becomes extremely patient.
I identify two conditions that help ameliorate these problems. These conditions are an incentive compatibility condition on the effective rate of future discounting: the long-run player wishes to play the Stackelberg action if recommended to do so, supposing maximal variation in continuation payoffs after; and a credible belief condition, represented by the Markov refinement: short-run players believe the long-run player has an incentive to build a reputation when those incentives are present. Applied to a global games setting, these conditions yield new predictions about the survival of equilibria relative to the risk dominant threshold in a dynamic regime change game.
There are several fruitful directions for future analysis. First, understanding the limits of endogenous exit when players’ actions are imperfectly monitored is a natural next step. The methods developed in this paper need not apply because reputation formation and incentive provision are much more subtle with imperfect monitoring, and the standard law-of-large-numbers intuitions do not apply with endogenous exit. Second, understanding the case with multiple commitment types is also promising, away from sequential equilibria—here, the set of payoffs may be pathological because of continuation payoffs at off-path deviations (i.e. taking a committed action that is not on the path of play). Understanding the set of Nash equilibria in this game and their payoff bounds is potentially promising. Finally, there is scope to apply the machinery in this paper to several other settings. For example, in the repeated trust game, my findings can give predictions about when the long-run player can teach short-run players to trust them even when a lack of trust might prematurely end the game. Alternatively, Theorem 1 can be used to give conditions under which Markov equilibria have reputation foundations in stopping time games where short-run or myopic players can stop at any time (i.e. adapted versions of Chassang (2010)). Finally, outside of Markov predictions, these results may give new interpretations of “bad reputation” style results in the spirit of Ely and Valimaki (2003) and Ely et al. (2008).
Appendix: Omitted Proofs
PROOF OF PROPOSITION 1
Proof.
Some preliminaries. First, $\underline{V}_{0}$ must be the long-run player’s lowest possible equilibrium payoff as 0 is dominant in the stage game. Similarly, $V^{*}$ must be their highest equilibrium payoff because short-run players must myopically best reply. Next, since $\mu_{0}(h^{t})=1$, we consider assessments where $\mu(h^{t})(\omega_{R})=1$ always and find subgame perfect equilibrium. Finally, throughout the proof, assume players have access to a public randomization device so that the payoff set is convex.
It is clear that $\underline{V}_{0}$ is attainable by an equilibrium where the lowest stage-game Nash equilibrium is played at every history. Suppose now the second condition is satisfied. Consider the equilibrium with the following strategies.
  1. (1)
    If $h^{t}=\{c^{*}\}_{s=0}^{t-1}$, $\sigma_{L}(h^{t})=c^{*}$, $\{\sigma_{i}(h^{t})\}_{i\in[N]}=a^{*}\in\mathcal{N}(c^{*})$.
  2. (2)
    At any other history, $\sigma_{L}^{*}(h^{t})=0$ and $\{\sigma_{i}(h^{t})\}_{i\in[N]}=\underline{a}\in\mathcal{N}(0)$, with $\underline{a}$ chosen to attain the lowest Nash equilibrium payoff $\underline{V}_{0}$.
By construction short-run players are myopically best replying. Clearly at histories where $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$ the long-run player’s strategy is sequentially rational. At a history where only $c^{*}$ has been played, Assumption 1 implies the long-run player’s most profitable deviation is to play 0; this is unprofitable so long as
$f(a^{*},c^{*})[v(a^{*},c^{*})+V^{*}]\geq f(a^{*},0)[v(a^{*},0)+\underline{V}_{0}].$
Divide through by $f(a^{*},0)$ and collect terms on the left hand side to conclude that the above expression is equivalent to
$(V^{*})L(a^{*},c^{*})-\underline{V}_{0}\geq v(a^{*},0)-v(a^{*},c^{*})L(a^{*},c^{*})$
as desired. Thus, $\underline{U},\overline{U}]=[\underline{V}_{0},V^{*}]$.
The converse. Suppose the second condition fails. Some algebra implies it must be that for any possible continuation payoffs $V(c^{*}),V(0)$ for playing $c^{*}$ and 0 respectively, we have
$f(a^{*},c^{*})[v(a^{*},c^{*})+V(c^{*})]<f(a^{*},0)[v(a^{*},0)+V(0)]$
and so it is never possible to induce the long-run player to play $c^{*}$ on path. Thus the long-run player’s largest payoff is bounded away from $\underline{V}_{*}$. Note this argument holds even for the discounted repeated game, for any discount rate. ∎
PROOF OF PROPOSITION 2
Proof.
First, note that if $\mu_{0}(\omega_{R})=1$, then it must be that in every sequential equilibrium $\mu(h^{t})(\omega_{R})=1$ as well (as the agent has a degenerate prior belief). Now fix any history $h^{t}$; at any successor history $h^{s},h^{s^{\prime}}\succ h^{t}$, it must be that $\sigma(h^{s})=\sigma(h^{s^{\prime}})$, so in particular future play cannot vary based on play today. Thus, the long-run player’s action must be myopically optimal, and hence $\sigma_{L}(h^{t})=0$ at any history $h^{s}$. Short-run players must play the same myopic best response at each history, so in equilibrium $\{\sigma_{i}(h^{t})\}_{i\in[N]}\in\mathcal{N}(0)$ always. Hence all equilibria are repeated static Nash. Public randomization convexifies the payoff set. ∎
PROOF OF LEMMA 1
Proof.
Let $(\sigma,\mu)$ be a sequential equilibrium and let $(\sigma^{n},\mu^{n})$ be the fully supported strategy which converges to it. Now fix any history $h^{t}\neq\{c^{*}\}_{s=0}^{t-1}$ noting it is supported by $\mathbb{P}^{\sigma^{n}}$ for all $n$. It must be that $\mu^{n}(h^{t})(\omega_{R})=1$, since no committed type could have produced this history but there is positive probability the rational type did. This implies $\lim\limits_{n\to\infty}\mu^{n}(h^{t})(\omega_{R})=\mu(h^{t})(\omega_{R})=1$, finishing the proof. ∎
PROOF OF LEMMA 2
Proof.
Let $(\sigma,\mu)$ be Markov perfect. By Proposition 2, any continuation strategy at histories $h^{t}$ where $\mu(h^{t})(\omega_{R})$ must be a repeated static Nash profile, so in particular $\sigma_{L}(h^{t})=0$.
Consider now a history $h^{t}$ where $\mu(h^{t})(\omega_{R})<1$ (noting these are all on path), and suppose the lemma is false. Then there exists some on-path continuation history $h^{t}$ where at $h^{t}\succ h^{t-1}$, $h^{t-1}=\{c^{*}\}_{t=0}^{t-2}$ but $h_{t}=c\not\in\{0,c^{*}\}$. Thus, $\mu(h^{t})(\omega_{R})=1$ and the long-run player’s continuation value at $h^{t}$ must be the same as their continuation value at $\{h^{t-1},0\}$. But then playing 0 instead of $c$ for sure is a strictly profitable deviation by Assumption 1, contradicting sequential rationality of $(\sigma,\mu)$. ∎
PROOF OF LEMMA 3
Proof.
Suppose not. Let $(\sigma,\mu)$ be an MPE and let $h^{t}$ be a history at which the long-run player mixes between 0 and $c^{*}$. By the argument in Lemmas 1 and 2, this can only possibly occur on a history $h^{t}=\{c^{*}\}_{s=0}^{t-1}$ for some $t$; call such a history the time-$t$ clean history. Define
$\mathcal{T}=\{t:h^{t}\text{ is clean and }\sigma_{L}(h^{t})\text{ mixes.}\}$
First, $|\mathcal{T}|$ cannot be infinite. Suppose not, so there is a sequence of times $\{t_{n}\}_{n=1}^{\infty}\in\mathcal{T}$. Note that when the long-run player mixes, both (1) $h^{t_{n}}\in\mathbb{P}^{\sigma}$ where $h^{t_{n}}$ is clean, and (2) the long-run player’s action is statistically identifiable compared to the committed type. This implies by a concentration of measure argument (see Mailath and Samuelson (2006) Chapter 15) that $\lim\limits_{n\to\infty}\mu(h^{t_{n}})(\omega_{c^{*}})=1$. Because short-run player actions are upper hemi-continuous in $\mu(h^{t_{n}})$, there must exist $N$ sufficiently large that for all $n\geq N$, $\mathcal{V}(h^{t_{n}}|\sigma)\geq\mathcal{V}(c^{*},h^{t_{n}})\geq V_{*}-\varepsilon$ for some $\varepsilon>0$, where $\mathcal{V}(c^{*},h^{t})$ is the long-run player’s continuation history from simply always taking $c$ in the future and $\mathcal{V}(h^{t}|\sigma)$ is their continuation value from $\sigma$. Yet mixing implies the long-run player must be indifferent between taking 0 (and obtaining at most $\overline{V}_{0}$ hereafter, since they have revealed their rationality) and taking $c^{*}$ (and obtaining some other future continuation value). In particular, for $N$ sufficiently large and $a=\{\sigma_{i}(h^{t_{N}})\}$, we have
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]=f(a,c^{*})[v(a,c^{*})+V(\{h^{t_{n}},c^{*}\}|\sigma)].$
Yet we also know from Proposition 2 that
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]\leq f(a,0)[v(a,0)+\overline{V}_{0}]$
and from our observation above that
$f(a,c^{*})[v(a,c^{*})+V(\{h^{t_{n}},c^{*}\}|\sigma)]\geq f(a,c^{*})[v(a,c^{*})+(V_{*}-\varepsilon)].$
Yet the assumption the long-run player satisfies $c^{*}$ IC implies
$f(a,0)[v(a,0)+\overline{V}_{0}]<f(a,c^{*})[v(a,c^{*})+V_{*}]$
for $a$ since $a\in\mathcal{N}(\alpha c)$. Taking everything together and using $c^{*}$-incentive compatibility implies for all $N$ sufficiently large that
$f(a,0)[v(a,0)+\mathcal{V}(\{h^{t},0\}|\sigma)]<f(a,c^{*})[v(a,c^{*})+V(\{h^{t_{n}},c^{*}\}|\sigma)]$
contradicting the indifference condition which is necessary to mix.
Thus, $|\mathcal{T}|$ is finite. Let $T>0$ be the last time at which the long-run player mixes at a clean history. This implies after $T$, the long-run player attains at least $V_{*}$ when they play 0. Thus, for $a=\{\sigma_{i}(h^{T})\}\in\mathcal{N}(\alpha c)$ for some $\alpha\in[0,1]$, we have
$f(a,0)[v(a,0)+\underline{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V^{*}]$
This again contradicts $c^{*}$ IC. Hence the long-run player never mixes. ∎
PROOF OF THEOREM 1 (Converse)
Proof.
Suppose the condition fails. There are three cases. First, suppose the inequality fails strictly at $\alpha=0$. Then in particular
$f(a,0)[v(a,0)+\overline{V}_{0}]>f(a,c^{*})[v(a,c^{*})+V_{*}]$
for $a\in\mathcal{N}(0)$. This incentive compatibility constraint implies that the rational long-run player would always prefer to separate and play 0 (starting from any time-0 belief) if they are guaranteed $\overline{V}_{0}$ afterwards. In particular, it implies the following is a Markov equilibrium as $\mu_{0}(\omega_{R})\to 1$:
  1. (1)
    The rational long-run player plays 0 in every period.
  2. (2)
    Short-run players play $a_{*}=a^{*}(c^{*})$ at all $h^{t}$ where $\mu(h^{t})(\omega_{R})=0$, with $a_{*}$ chosen to deliver the long-run player expected continuation payoff $V_{*}$.
  3. (3)
    If $\mu(h^{t})(\omega_{R})=1$, play $a_{0}^{*}\in\mathcal{N}(0)$ chosen to deliver the long-run player an expected continuation payoff $\overline{V}_{0}$.
  4. (4)
    At histories where $\mu(h^{t})(\omega_{R})=\mu_{0}(\omega_{R})$, play any $a\in\mathcal{N}(\mu_{0}(\omega_{c^{*}})c^{*})$.
Second, suppose instead that the inequality fails for some $\alpha\in(0,1)$ but that it holds for $\alpha=1$. This implies first that it is possible to sustain the Markov equilibrium where the long-run player plays $c^{*}$ at all histories. Moreover, there must exist some $\overline{\alpha}\in(0,1)$ with $a\in\mathcal{N}(\alpha c^{*})$ and
$f(a,0)[v(a,0)+\overline{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V_{*}]$
by the intermediate value theorem, since the incentive compatibility constraint is continuous in $\alpha$. This implies that there exists a Markov equilibrium of the following form:
  1. (1)
    If $\mu(h^{t})(\omega_{R})\neq\{\mu_{0}(\omega_{R}),1\}$, then have $\sigma_{L}(h^{t})=c^{*}$ and $\{\sigma_{i}(h^{t})\}\in\mathcal{N}(c^{*})$ be chosen to give the long-run player $V_{*}$.
  2. (2)
    If $\mu(h^{t})(\omega_{R})$, play static Nash $\sigma(h^{t})$ giving the long-run player $\overline{V}_{0}$.
  3. (3)
    If $\mu(h^{t})(\omega_{R})=\mu_{0}(\omega_{R})$, have $a=\{\sigma_{i}\}\in\mathcal{N}(\alpha c*)$ chosen so that
    $f(a,0)[v(a,0)+\overline{V}_{0}]=f(a,c^{*})[v(a,c^{*})+V_{*}]$
    and $\sigma_{L}(h^{t})\in(0,1)$ chosen so that
    $\mu_{0}(\omega_{R})\sigma(h^{t})(c^{*})+(1-\mu_{0}(\omega_{R}))=\alpha$
    Noting this is well-defined so long as $\mu_{0}(\omega_{R})>1-\alpha$.
It is clear that at all histories, the short-run player is best replying, as is the long-run player at all histories outside of $\varnothing$. Moreover, at the first history $h^{t}=\varnothing$, by choice of $\alpha$, the long-run player is indifferent between $c^{*}$ and 0 and thus is willing to randomize given $a$. Hence this is a Markov perfect equilibrium. This gives the long-run player an equilibrium payoff of
$f(a,0)[v(a,0)+\overline{V}_{0}]<f(a_{*},0)[v(a_{*},0)[v(a_{*},0)+\overline{V}_{0}]]$
$\leq f(a_{*},c^{*})[v(a_{*},c^{*})+V_{*}]$
where the first inequality follows from monotonicity of the game and the second follows from the fact the incentive compatibility condition holds at $\alpha=1$ where $a_{*}=a^{*}(c^{*})$ is the largest element of $\mathcal{N}(c^{*})$, since again by the monotonicity assumption this is the choice which minimizes the stage-game payoff. This implies the result.
Finally, suppose the inequality fails when $\alpha=1$. Then there cannot be an equilibrium where the long-run player imitates $c^{*}$ for sure. The only other pure Markov equilibrium gives the long-run player their minimax payoff and implies the result. If instead the long-run player mixes in a Markov equilibrium at some on-path history, they secure (for some $a\in\mathcal{N}(\alpha c^{*})$ and some $\alpha\in(0,1)$) a payoff of at most
$f(a,c^{*})[v(a,c^{*})+V_{*}]<f(a^{*},c^{*})[v(a^{*},c^{*})+V_{*}]=V_{*}$
by monotonicity and hence must secure a payoff smaller than $V_{*}$. ∎
Proof of Corollary 1
Proof.
First, note that under these conditions, the incentive compatibility constraint in the folk theorem can be written as
$(V^{*})\mathcal{L}(c^{*})-\underline{V}_{0}\geq v(0)-v(c^{*})\mathcal{L}(c^{*})$
and $c^{*}$ incentive compatibility can be written as
$(V^{*})\mathcal{L}(c^{*})-\underline{V}_{0}>v(0)-v(c^{*})\mathcal{L}(c^{*})$
because dependence on $a$ is suppressed in the likelihood ratio and $|\mathcal{N}(0)|$ is a singleton so $\underline{V}_{0}=\overline{V}_{0}$. These are the same incentive compatibility constraints. Hence the two statements are equivalent.
That there is a unique Markov equilibrium when the IC constraint is satisfied then follows by noting that (1) there is a unique static Nash equilibrium (and hence a unique strategy at all histories in an MPE where $\mu(h^{t})(\omega_{R})=1$) and (2) that $\Omega=\{0,c^{*}\}$ contains only two elements, so any histories where $\mu(h^{t})(\omega_{R})<1$ must be clean with a fixed element. Thus the unique MPE is the one that plays $c^{*}\in\Omega$ at every history. ∎
PROOF OF PROPOSITION 3
Proof.
Fix $\varepsilon>0$ and preferences $(u,v)$. Fix some action $c$ and $a^{*}(c)$ be some best reply to it where $v(a^{*}(c),c)>0$, which exists by the nontriviality assumption. Set $f(a^{*}(c),c)=1-\eta,f(a^{*}(c),0)=\eta$. Let $\underline{a}=a^{*}(0)\in\mathcal{N}(0)$ be such that $(\underline{a},0)$ gives the long-run player payoff $\underline{V}_{0}$. Set moreover $f^{*}(\underline{a},0)=f^{*}(\underline{a},c)=\eta$. Note for $\eta$ sufficiently small, playing $c$ must be preferred to playing 0 since
$\frac{v(a^{*}(c),c)(1-\eta)}{1-(1-\eta)}\geq\frac{v(a,0)\eta}{1-\eta}$
for any $a\in\mathcal{N}(0)$. This in particular implies for all $\eta>0$, there is an optimal commitment posture $c^{*}\neq 0$ and in particular one which guarantees a payoff of at least $v(a^{*}(c),c)>0$ in each period conditional on surviving.
Let $c^{*}(\eta)$ be the optimal commitment posture given the exit function described above (and $V_{*}(\eta)$ its associated Stackelberg payoff). In this case, Proposition 1 implies the folk theorem is equivalent to
$f(a^{*}(c^{*}(\eta)),c^{*})[v(a^{*}(c^{*}(\eta)),c^{*})+V_{*}(\delta,\eta)]\geq f(a^{*},0)[v(a^{*}(0)+\underline{V}_{0}]$
The term inside the brackets on the right hand side is independent of $\eta$, and the left hand side by assumption is bounded from below by
$f(a^{*}(c),c)[v(a^{*}(c),c)+V_{*}(\eta)]$
as $\eta$ vanishes. But then our condition is implied by the requirement that
$\frac{f(a^{*}(c),c)}{f(a^{*},0)}[v(a^{*}(c),c)+V_{*}(\eta)]\geq[v(a^{*}(0)+\delta V_{*}]$
$\iff$ $\frac{1-\eta}{\eta}[v(a^{*}(c),c)+V_{*}(\eta)]\geq[v(a^{*}(0)+\underline{V}_{0}]$
which is clearly true as $\eta\to 0$. Hence under this survival probability function $f^{*}$, $c^{*}$ optimality is satisfied for all $\eta$ sufficiently small.
Fix now some $\eta$ sufficiently small for the folk theorem to hold (in particular so that $V^{*}$ can be sustained as an equilibrium) under $f$ and consider the following (non-Markov) strategy profile at time $t=0$:
  1. (1)
    Have the long-run player play 0 in the first period.
  2. (2)
    Have the short-run player’s play to $a^{*}(\mu_{0}(\omega_{c^{*}})c^{*}))$, i.e. best respond to their prior belief supposing the rational type plays 0.
  3. (3)
    Regardless of the long-run player’s action, given them a payoff of $V^{*}$ tomorrow.
Clearly $V^{*}$ is sustainable by the folk theorem tomorrow; and because the rational long-run player’s action today does not affect their future continuation payoff, they will play their dominant action 0. Finally, Condition (2) is exactly the condition for short-run players to myopically best reply to this strategy by the long-run player. As $\mu_{0}(\omega_{R})\to 1$, short-run players (by upper hemi-continuity) converge to an action in $\mathcal{N}(0)$. In particular, this implies $\underline{a}$ must be short-run players a (weakly) lower payoff than $\lim\limits_{\alpha\to 0}v(a^{*}(\alpha c^{*}),c^{*})$.
This equilibrium strategy profile gives the long-run player a payoff of exactly
$f(\underline{a},c)[v(\underline{a},0)+V_{*}]=\eta[v(\underline{a},0)+V_{*}].$
Choosing $\eta<\frac{\varepsilon}{v(\underline{a},0)+\delta V_{*}}$ then implies the argument.
Finally, the long-run player’s payoff can be arbitrarily high. For fixed $\eta$, set $f(a^{*}(c^{*}(\eta),c^{*}(\eta)))=1-\varsigma$. Then
$V^{*}\geq(1-\varsigma)\frac{v(a^{*}(c^{*}(\eta)),c^{*}(\eta))}{\varsigma}$
which diverges as $\varsigma\to\infty$. ∎
PROOF OF PROPOSITION 4
Proof.
The first equivalence. Rank $F\succ A$ and $In\succ Out$. Then $f(a,c)$ and $v(a,c)$ are strictly decreasing for each fixed $c$, and $\mathcal{N}(\alpha\text{Fight})$ is decreasing in $\alpha$, so the game is monotone. Thus the first statement follows from Theorem 1 once we verify $c^{*}$-incentive compatibility. First, note that
$V_{*}=\frac{2(1-\eta)}{1-(1-\eta)}\text{ and }\overline{V}_{0}=\frac{\varepsilon}{1-\varepsilon}.$
When $a=Out$, $L(a,c^{*})=1$; hence the condition boils down to
$\left(\frac{2(1-\eta)}{1-(1-\eta)}-\frac{\varepsilon}{1-\varepsilon}\right)>(2-2)$
which is clearly satisfied since $1-\eta>\varepsilon$. When instead $a=In$, the condition becomes
$\left(\frac{2(1-\eta)}{1-(1-\eta)}\frac{\varphi}{\varepsilon}-\frac{\varepsilon}{1-\varepsilon}\right)>1$
Clearly the expression on the right hand side is decreasing in $\varepsilon$ and satisfied as $\varepsilon\to 0$ for fixed $\varphi$, and so in particular the intermediate value theorem furnishes the necessary $\varepsilon$ (possibly $\varepsilon=1$) for the result to hold.
The second part of the result follows by constructing the type of equilibrium in Proposition 3, taking $\varepsilon\to 0$ and noting that the long-run player will not fight in the first period because the “exit probability” is sunk. ∎
PROOF OF PROPOSITION 5
Proof.
Fix a value $c$, and note the induced stage-game is a finite monotone (decreasing) supermodular game with dominance regions $\infty,0)$ and $(1,\infty]$. This implies by the proof of Theorem 1 of Frankel et al. (2003) that for all $\varsigma$ sufficiently small, $s(x,c)=\mathbf{1}\left\{x<x^{*}(c)\right\}$ is the unique equilibrium strategy profile.
Next, let $U_{i}(c,x,x^{*})$ be the interim utility differential between taking $a=1$ and $a=0$ for an agent observing signal $x$ when players use a cutoff strategy $x^{*}$ and the cost of failed revolution is $c$. In equilibrium, it must be that $U_{i}(c,x^{*}(c),x^{*}(c))=0$. To ensure that $x^{*}(c)$ is differentiable and strictly increasing, we appeal to the implicit function theorem; some algebra gives that $x^{*}(c)$ is implicitly defined as the solution to the equation
$U_{i}(c,x,x)=(1+c)F_{\theta}(F_{x}(x|x)|x)-c=0.$
We have that
$\frac{\partial U_{i}}{\partial c}=F_{\theta}(F_{x}(x|x)|x)-1\text{ and }\frac{\partial U_{i}}{\partial x}=(1+c)\left(F_{\theta}(F_{x}(x|x)|x)\cdot f_{x}(x|x)\cdot\frac{\partial}{\partial 2}F_{\theta}(F_{x}(x|x)|x)\cdot\frac{\partial}{\partial 2}F_{x}(x|x)\right)$
where $\frac{\partial}{\partial 2}$ refers to the derivative in the conditioning argument. Note the derivative in $c$ is strictly negative as $P_{\theta}(\cdot)<1$ always; meanwhile, the second term is strictly negative by Assumption 2. Thus, the implicit function theorem implies
$\frac{\partial x^{*}(c)}{\partial c}=-\frac{\partial U_{i}}{\partial c}\left(\frac{\partial U_{i}}{\partial x}\right)^{-1}<0,$
as desired. The last statement follows similarly. Regime change only occurs if $F_{x}(x^{*}(c)|\theta)>\theta$, so $\theta^{*}(c)$ is implicitly defined by the equation $F_{x}(x^{*}(c)|\theta)-\theta=0$. Differentiating yields
$\frac{d}{d\theta}(F_{x}(x^{*}(c)|\theta)-\theta)=\frac{\partial}{\partial 2}F_{x}(x^{*}(c)|\theta)-1<0$
as $\frac{\partial}{\partial 2}F_{x}(x|\theta)<0$ for all $x$; this then implies $\theta^{*}(c)$ is globally $\mathcal{C}^{1}$. Applying the implicit function theorem again implies its derivatives is given by
$\frac{\partial\theta^{*}(c)}{\partial c}=-\left(f_{x}(x^{*}(c)|\theta)\frac{\partial x^{*}(c)}{\partial c}\right)\left(\frac{\partial}{\partial 2}F_{x}(x^{*}(c)|\theta)-1\right)^{-1}<0,$
and so higher costs $c$ decrease the region of states where attacks are successful. ∎
PROOF OF THEOREM 2
Proof.
Fix some distribution $P_{\theta}$. Note
$\lim\limits_{\varsigma\to 0}F_{x}\left(\frac{1}{1+c^{*}}\bigg|x\right)\to\begin{cases}0\text{ if }x<\frac{1}{1+c^{*}}\\ 1\text{ if }x>\frac{1}{1+c^{*}}\end{cases}$
because $F_{x}(\cdot|\theta)$ converges to $\delta_{\theta}$ as $\varsigma\to 0$ (people become extremely sure about the state). Since $\theta^{*}(c)$ is continuous and zeros $F_{x}\left(\frac{1}{1+c^{*}}|\theta\right)-\theta$, this implies $\theta^{*}(c)=\frac{1}{1+c^{*}}$.
The first statement. Suppose $\frac{1}{1+c^{*}}<\mu$ and let $\eta=\mu-\frac{1}{1+c^{*}}$, noting this gap is independent of $\tau$. We want to prove
$V^{*}-\underline{V}_{0}>\kappa(c^{*})$
in the right order of limits. Set $\overline{P}_{\theta}(x)=1-P_{\theta}(\theta^{*}(x))$ to be the probability of survival given a commitment posture. Some algebra implies that incentive compatibility is equivalent to the requirement
$\overline{P}_{\theta}(c^{*})-\overline{P}_{\theta}(0)\geq\kappa(c^{*})[1-\overline{P}_{\theta}(0)].$
Note from here that
$\overline{P}_{\theta}(c^{*})=1-P_{\theta}(\theta^{*}(c^{*}))>1-P_{\theta}(\mu-\eta)\geq 1-\frac{\text{var}(\theta)}{\eta^{2}}>1-\frac{\overline{\tau}}{\eta}.$
The middle inequality uses Chebychev’s inequality. Thus $\overline{P}_{\theta}(c^{*})\to 1$ as $\overline{\tau}\to 0$; similarly $\overline{P}_{\theta}(0)\to 0$ as $\overline{\tau}\to 0$; together this implies the incentive constraint holds. Theorem 1 and Corollary 1 imply the regime can sustain $c^{*}$ and some best response to $c^{*}$ in the unique Markov equilibrium, as desired. Finally because $\overline{P}_{\theta}(c^{*})=f(c^{*})$, the probability of exit vanishes as $\overline{\tau}\to 1$. This implies both parts of the first statement, noting $\overline{\varsigma}$ can depend on the choice of $\tau$.
Now consider the case where $\frac{1}{1+c^{*}}>\mu$, now setting $\eta=\frac{1}{1+c^{*}}-\mu>0$. Note a necessary condition for incentive constraints to hold is for
$\frac{\overline{P}_{\theta}(c^{*})-\overline{P}_{\theta}(0)}{1-\overline{P}_{\theta}(0)}>\kappa(c^{*})$
As $\bar{\tau}$ vanishes, one has that
$1-P_{\theta}(\theta^{*}(\sigma^{*}(h^{t}))\leq 1-P_{\theta}(\theta^{*}(c^{*}))<1-P_{\theta}(\mu+\eta)\leq\frac{\text{var}(\theta)}{\eta^{2}}<\frac{\overline{\tau}}{\eta^{2}}$
again using Chebychev’s inequality. Since $\theta^{*}(c)<\theta^{*}(0)$, we conclude that
$\frac{\overline{P}_{\theta}(c^{*})-\overline{P}_{\theta}(0)}{1-\overline{P}_{\theta}(0)}\leq\frac{\frac{\bar{\tau}}{\eta^{2}}-\bar{P}_{\theta}(0)}{1-\bar{P}_{\theta}(0)}.$
Hence this term is bounded from above by $\kappa(c^{*})$ for all sufficiently small $\tau$. Thus, $c^{*}$-incentive compatibility must fail strictly, so by the converse Theorem 1 in the unique Markov equilibrium the regime plays 0 in every period. Since $\overline{P}_{\theta}(0)$ vanishes, $\underline{U}_{0}<\varepsilon$. As this was the unique Markov equilibrium, $\underline{U}^{M}<\varepsilon$ as well. This finishes both directions of the proof. ∎
SUPPLEMENTAL APPENDIX
This appendix considers the discounted repeated game at some discount rate $\delta$. Section B.1 discusses some problems with normalized repeated games, interpreted as the average payoffs across periods. Section B.2 derives a novel comparative statics result on the set of commitment postures as the discount rate changes, and interprets it in the context of applications. Section B.3 sets the main result in the language of the discounted game.
B.1. Non-normalized payoffs
Intuitively, without exit, the stream of normalized discounted payoffs converges to the average payoff as $\delta\to 1$. However, when the game has endogenous exit, so long as exit occurs with high enough probability, the average stream of payoffs will be zero for “most” periods (i.e. if the expected exit time is finite), so normalized payoffs will vanish as players get extremely patient. The lemma formalizes this observation.
Lemma B.1.
Let $\{f_{t}\}\subset[0,1]$ be a sequence of exit probabilities. If $\liminf_{t\to\infty}f_{t}<1$, then
$\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=0}^{\infty}v_{t}\prod_{s=0}^{t}f_{s}=0.$
for any feasible payoffs $v_{t}$.
Proof.
Let $\{\overline{f}_{\tilde{t}}\}\subset\{f_{t}\}$ be a subsequence such that $\lim\limits_{\tilde{t}\to\infty}\overline{f}_{\tilde{t}}=\overline{f}<1-\eta$, for some $\eta>0$ fixed; moreover, by truncating finitely many terms at most, suppose $\overline{f}_{\tilde{t}}<1-\eta$ for all $s\in\mathbb{N}$. Now for any $T\in\mathbb{N}$, define the number of times that $f_{t}<1-\eta$ to be $s(T)=\#\{s:f_{s}<1-\eta\}$. By assumption of $\eta$, we know $\lim\limits_{T\to\infty}s(T)=\infty$. We now have that $\prod_{s=0}^{\infty}f_{t}<(1-\eta)^{s(T)}$. Letting $\overline{v}$ be chosen so that $|v_{t}|\leq\overline{v}$ for all $t$, we have
$0$ $\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}v_{t}\prod_{s=0}^{\infty}f_{s}\right|\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{\infty}\overline{v}(1-\eta)^{s(t)}\right|$
$\leq\lim\limits_{\delta\to 1}\left|(1-\delta)\sum_{t=0}^{T}\overline{v}(1-\eta)^{s(t)}\right|+\left|(1-\delta)(1-\eta)^{s(T)}\sum_{t=T+1}^{\infty}\overline{v}\right|$
$\leq(1-\eta)^{s(T)}\lim\limits_{\delta\to 1}(1-\delta)\sum_{t=T+1}^{\infty}\overline{v}\leq(1-\eta)^{s(T)}\overline{v}.$
Since this holds for all $T\in\mathbb{N}$, taking $T\to\infty$ (and hence $s(T)\to\infty$) gives the argument. ∎
Lemma B.1 implies that it is not fruitful to compare discounted average payoffs in the patient limit, as they will often be zero even when payoffs are meaningfully different.
There are two potential ways to ameliorate this problem. The first is to normalize by the effective discount rate, $f(a,c)\delta$, in each period. However, such a normalization is infeasible and difficult to interpret: $f(a,c)$ is an endogenous object, and thus one should naturally expect that it vary the long-run players’ payoffs. Moreover, except in special cases, the stream of exit probabilities $\{f(a_{t},c_{t})\}_{t=0}^{\infty}$ need not be time-invariant, making it hard to understand what the “average effective discount rate” should be. The second way is to simply work directly with the undiscounted payoffs $\mathcal{V}(\{a_{t},c_{t}\})$; I take this approach in this paper.
This choice, however, leads itself to a novel technical challenge: without normalizing payoffs by the discount rate, deviations today may remain first order (in particular, they are not order $(1-\delta)$ compared to an order $\delta$ variation in the continuation payoff, and are instead order $1$ versus order $\delta$ variation in the continuation payoff). This makes it much harder to discipline the long-run player from deviating to play non-Stackelberg actions, since deviations from potential equilibrium profiles will always remain tempting. Theorem 1 can thus be interpreted as providing a simple economic condition under which the long-run player can successfully establish a (Markov) reputation, despite this hurdle.
B.2. Commitment Postures
How does the discount rate affect the long-run player’s Stackelberg payoff? Define
$c^{*}=\operatorname*{\arg\!\max}_{c\in C}\min_{a\in\mathcal{N}(c)}\sum_{t=0}^{\infty}f(a,c)^{t+1}v(a,c)$
to be the set of (lower) Stackelberg actions for the long-run player when they discount the future at rate $\delta$. Note the long-run player trades off between a higher survival probability at the (potential) cost of a lower stage-game payoff in each period. My next result characterizes exactly how the long-run player resolves this tradeoff as they get more patient: regardless of the speed at which $v(a,c)$ might decline in $c$, they will always favor more aggressive actions that increase their survival probability as they get more patient. To do so, I make one more assumption on the survival probability.
Assumption 3.
$f^{*}(c)=\min_{a\in\mathcal{N}(c)}f(a,c)$ is nondecreasing in $c$.
Proposition B.1 (Patience Effect).
Suppose Assumption 3 is satisfied. $c^{*}$ is nonempty, upper hemicontinuous, and increasing (in the strong set order) in $\delta$.
Proof.
We first claim $f^{*}(c)$ is continuous, since for any sequence $c_{n}\to c$,
$\lim\limits_{n\to\infty}f^{*}(c_{n})=\lim\limits_{n\to\infty}\frac{v(a^{*}(c_{n}),c_{n})f(a^{*}(c_{n}),c_{n})}{1-f(a^{*}(c_{n}),c_{n})}=\frac{v(a^{*}(c),c)f(a^{*}(c),c)}{1-f(a^{*}(c),c)}.$
The first inequality is by definition, and the second by upper hemi-continuity of the Nash correspondence. Given continuity of $f^{*}(c)$, Berge’s theorem of the maximum applied to the optimizing set $c^{*}=\max_{c\in C}f^{*}(c)$ implies $c^{*}$ exists and is upper hemi-continuous. Monotonicity now relies on the following lemma [^18] Such a lemma appears elementary, but to the best of my knowledge there is not a prior reference. Note the lemma establishes a general comparative statics result for discounted streams of payoffs that cross zero once from below..
Lemma B.2.
Let $\{b_{t}\}_{t\in\mathbb{N}}\subset\mathbb{R}$ be a bounded sequence of real numbers, not identically zero, such that $b_{t}<0$ if and only if $t<T$ for some $T\in\mathbb{Z}_{+}$. Then for any $\delta_{1}<\delta_{2}$,
$\sum_{t=0}^{\infty}\delta_{1}^{t}b_{t}\geq 0\implies\sum_{t=0}^{\infty}\delta_{2}^{t}b_{t}>0.$
Proof.
Note
$\sum_{t=0}^{\infty}\delta_{1}^{t}b_{t}=\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}\geq 0\iff\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}$
As a function of $\delta$, this series is absolutely convergent for all $\delta\in(0,1)$ and the partial sums converge uniformly on this interval so we can differentiate under the summation: thus,
$\frac{d}{d\delta}\sum_{t=0}^{\infty}\delta_{1}^{t-T}b_{t}=\sum_{t=0}^{\infty}(t-T)\delta_{1}^{t-T-1}b_{t}<\infty$
where we use the fact $b_{t}$ is bounded to guarantee the sum is well-defined. Note each term in this summation is nonnegative. If $t<T$, then tT<0t-T<0 and $b_{t}<0$, so $(t-T)\delta^{t-T-1}b_{t}>0$, while if $t>T$, then $(t-T)\delta^{t-T-1}\geq 0$ and $b_{t}\geq 0$ as well. Moreover, since $\{b_{t}\}$ is not identically zero, at least one term in the sum above must be strictly positive, and hence the derivative is strictly positive. We thus have that
$\sum_{t=0}^{\infty}\delta_{2}^{t-T}b_{t}>0\implies\delta_{2}^{T}\sum_{t=0}^{\infty}\delta_{2}^{t-T}b_{t}=\sum_{t=0}^{\infty}\delta_{2}^{t}b_{t}>0$
where we use the fact $\delta_{2}>\delta_{1}$, and finally multiply by $\delta_{2}^{T}>0$. ∎
Fix any $\delta$ and $c^{\prime}>c$; by Assumption 3, it must be that for all $(c^{\prime},c)$, $f^{*}(c^{\prime}),f^{*}(c)>0$. Consider the difference
$\sum_{t=0}^{\infty}[v(a^{*}(c^{\prime}),c^{\prime})f^{*}(c^{\prime})^{t+1}-v(a^{*}(c),c)f^{*}(c)^{t+1}]$
where I abuse notation to and have $a^{*}(\cdot)$ be any selection from the minimizing correspondence. Note that if $b_{t}$ is defined as
$b_{t}=v(a^{*}(c^{\prime}),c^{\prime})f^{*}(c^{\prime})^{t+1}-v(a^{*}(c),c)f^{*}(c)^{t+1}$
then
$b_{t}\geq 0\iff\frac{v(a^{*}(c^{\prime}),c^{\prime})}{v(a^{*}(c),c)}\geq\left(\frac{f^{*}(c)}{f^{*}(c^{\prime})}\right)^{t+1}.$
The left hand side is strictly positive, while the right hand side is positive, in $(0,1)$ (by Assumption 3), and strictly decreasing in $t$. Thus, there exists a unique $T$ (possibly 0) at which the inequality holds for all $t\geq T$. Lemma B.2 then implies, for any $c^{\prime}>c$, $\sum_{t=0}^{\infty}b_{t}$ is increasing in $\delta$, and hence the objective function $\sum_{t=0}^{\infty}v(a^{*}(c),c)$ is single-crossing in $(\delta,c)$. The result then follows by the Milgrom and Shannon (1994) comparative statics theorem. ∎
Proposition B.1 applies regardless of the speed or curvature of $v$ (in particular, it holds even if $v$ declines very fast relative to $c$), though this comes at the cost of a weak comparative static (every selection from $c^{*}$ is nondecreasing in $c$). Moreover, it has economic implications in the context of the applications—a regime seeking to fend off a coordinated attack will invest more in survival when they are more patient; longer-lived firms care more about ensures entrants are deterred; patient motorcyclists have a greater incentive to take “separating” actions that prove they are not commitment types, in order to deter exit.
B.3. The Main Result
That $c^{*}$ varies with the discount rate produces a new order of limits challenge in trying to prove some version of a commitment payoff theorem in the discounted repeated game. In particular, since the long-run player’s Stackelberg action depends on the discount rate, it is not clear that there is a finite set $\Omega$ such that $\Omega\cap c^{*}$ is nonempty for all $\delta$ close to $1$. Consequently, for any fixed set of commitment types, it may be infeasible to approximate a patient committed sender.
To obviate this problem, I index and simplify the set of commitment postures, $\Omega=\{\omega_{R},\omega_{c^{*}}\}$ (where I abuse notation and use $c^{*}$ to refer to a single selection from the set of commitment types). Moreover, as before, I set $\underline{U}$ and $\overline{U}$ as the infimum and supremum of the long-run player’s best and worst sequential equilibrium payoff, though now I implicitly fix the set of commitment types $\Omega=\Omega$, though I still suppress dependence on the specific sequence of prior beliefs $\mu_{0}\in\Delta(\Omega)$ (which itself my also change in $\delta$). Give this notation, the limit infimum (supremum) $\liminf_{\delta\to 1}\underline{U}$ (resp. $\limsup_{\delta\to 1}$) should be interpreted as the limit infimum (supremum) of the sequence of lower equilibrium payoffs in the repeated game with discount rate $\delta$, commitment types $\Omega$, and some fixed sequence of prior beliefs $\mu_{0}\in\Delta(\Omega)^{o}$. This interpretation implies that statements of the form $\limsup_{\delta\to 1}|V_{*}-\underline{U}|=0$ can be read as “The long-run player attains almost their Stackelberg payoff so long as there is positive probability players believe they are committed to their Stackelberg action at their level of patience”.
Under this interpretation, the following definition and result are immediate applications of the undiscounted game, by continuity of the payoff parameters in the discount rate.
Definition B.1.
Say the long-run player asymptotically satisfies (weak) $c^{*}$ incentive compatibility if there exists $\overline{\delta}<1$ such that for all $\delta>\overline{\delta}$ and each $a\in\mathcal{N}(\alpha c)$ for $\alpha\in[0,1]$,
$\delta\left(V_{*}(\delta)L(a,c^{*}(\delta))-\overline{V}_{0}(\delta)\right)>(\geq)\left(v(a,0)-v(a,c^{*}(\delta))L(a,c^{*}(\delta))\right)$
Theorem 3.
Let $\Omega(\delta)=\{0,c^{*}(\delta)\}$. If the long-run player asymptotically satisfies $c^{*}$ incentive compatibility, then for all $\mu_{0}$ with $\mu_{0}(\omega_{R})<1$,
$\limsup_{\delta\to 1}|\underline{U}^{M}(\delta)-V_{*}(\delta)|=0.$
Conversely, if $(f,v,u)$ is monotone and the long-run player does not asymptotically satisfy weak $c^{*}$ incentive compatibility, then for all $\delta<1$,
$\limsup_{(\delta,\mu_{0}(\omega_{R}))\to(1,1)}|\underline{U}^{M}(\delta)-V_{*}(\delta)|>0.$